### v1

# Cascade-driven opinion dynamics on social networks

Abstract
Online social networks (OSNs) have transformed the way individuals fulfill their social needs and consume information. As OSNs become increasingly prominent sources for news dissemination, individuals often encounter content that influences their opinions through both direct interactions and broader network dynamics. In this paper, we propose the Friedkin-Johnsen on Cascade (FJC) model, which is, to the best of our knowledge, is the first attempt to integrate information cascades and opinion dynamics, specifically using the very popular Friedkin-Johnsen model. Our model, validated over real social cascades, highlights how the convergence of socialization and sharing news on these platforms can disrupt opinion evolution dynamics typically observed in offline settings. Our findings demonstrate that these cascades can amplify the influence of central opinion leaders, making them more resistant to divergent viewpoints, even when challenged by a critical mass of dissenting opinions. This research underscores the importance of understanding the interplay between social dynamics and information flow in shaping public discourse in the digital age.
Index Terms:
Information cascades, opinion dynamics, polarization, Friedkin-Johnsen model, viral content
I Introduction
Socialization represents a fundamental aspect of human behavior, with empirical evidence consistently demonstrating the correlation between human health and social engagements [1, 2, 3], underscoring how isolation has serious consequences on both physical and mental health conditions [4, 5, 6], and could potentially lead to premature mortality [7]. The advent of online social networks (OSNs) has significantly reshaped how individuals fulfill their social needs, offering unprecedented opportunities to connect with others. Additionally, OSNs have emerged as prominent sources of real-time news dissemination, with a substantial proportion of the population relying on these platforms for information [8]. Political figures, journalists, and ordinary users frequently utilize OSNs to disseminate announcements and opinions, rendering them integral to contemporary information consumption patterns. Moreover, media organizations have implemented strategies to encourage the sharing of their content on OSNs, facilitating the seamless integration of news into users’ social interactions. Therefore, during our social activities online, we can also be exposed to the news even incidentally. However, this convergence of socialization and news consumption on OSNs has raised concerns regarding the formation of filter bubbles and echo chambers, wherein individuals are exposed primarily to content that reinforces their existing beliefs. Such environments pose risks of fostering biased perspectives and perpetuating misinformation, particularly with the proliferation of fake news campaigns designed to manipulate public opinion [9]. These dynamics underscore the complex interplay between socialization, news consumption, and information dissemination in the digital era, warranting further empirical investigation and theoretical exploration within interdisciplinary frameworks.
The dynamics of opinion formation represent a multifaceted research area situated at the intersection of various disciplines, including sociology, economics, psychology, control theory, computational sociology, and physics. Numerous models have been proposed to describe the mechanisms underlying opinion dynamics, aiming to capture observed theoretical and empirical phenomena. These models often account for factors such as individuals’ inclination to reduce cognitive dissonance with close acquaintances or their resistance to relinquishing deeply held biases. Among the prominent models in this domain is the Friedkin-Johnsen (FJ) model [10] belonging to the category of averaging models [11, 12, 10], according to which people iteratively adjust their opinions towards the average opinion of their neighbours, with the strength of social ties weighting the influence of each neighbor’s opinion. Additionally, the FJ model incorporates a parameter representing individuals’ open-mindedness, referred to as susceptibility, which modulates the extent to which individuals revise their opinions. The FJ model has gained popularity due to its mathematical tractability and its empirical validation on real social networks. It has demonstrated the ability to replicate various observed mechanisms in social networks, including the phenomenon of wisdom of the crowd [13] and the emergence of polarization [14] and recently, it has also been proposed as the building block to model multilateral international negotiation processes [15].
The FJ model has undergone rigorous validation in offline scenarios, where opinions evolve reflecting real-world interactions and debates on various topics. However, the dynamics change when socializing online, as we are exposed to friends’ opinions not through direct interaction but via their public activity on social media platforms, such as posting or commenting on news articles. Consequently, we may frequently encounter cognitive dissonance toward opinions that are expressed more frequently online. These online opinions often garner larger-scale attention and spread more intensely than their offline counterparts, potentially heightening the impact of social opinion exchange in digital social environments. These opinions are conveyed via textual posts, resharing of news items, or images. The dynamics of how content items are shared and spread online are studied under the name of information cascades [16].
Opinion dynamics and information cascades operate at distinct granularity and temporal scales. Information cascades focus on how individual pieces of information (e.g., a post, an image) are shared and propagate within the social network, while opinion dynamics concern the evolution of opinions through interactions among individuals. Both processes exhibit emergent behaviors, transitioning from local interactions to global effects such as content virality and polarization. In opinion dynamics, the specific social interactions leading to opinion change remain abstract and unspecified. Conversely, information cascades emphasize the sharing mechanism of content without addressing the effects of such sharing. However, given that reposting online content received from friends is a common online interaction, it is reasonable to expect an intertwining of opinion dynamics and information cascades. Despite this, the two processes have thus far been treated separately in the literature. In this work, we set out to fill this gap and establish an integrated model whereby the microscopic act of content sharing bears an effect in the macroscopic phenomenon of opinion evolution in online social networks.
To address this issue, we propose a novel model that integrates the FJ model with information cascades. Specifically, we hypothesize that opinions on a selected topic are expressed following the dissemination flow of online posts. Initially shared by certain sources, the post propagates as a cascade to nodes within their immediate and subsequent neighborhoods until it loses interest or reaches all nodes. Each shared post is assumed to convey the opinion of the node on the selected topic, e.g., expressed in text as feedback to a shared news article, thereby triggering cognitive responses from neighbors. The reception of a post from a neighbor, triggering an opinion update on the receiving node.
Furthermore, building upon our previous work [14], we analyze the impact of the cascade on polarization, evaluating the polarization values on the opinion vectors that maximize polarization of the classical FJ model. We tested the model over synthetic and real social networks. Our research reveals that the cascade-driven FJ model amplifies the influence of network topology compared to the classical FJ model. This empowers central nodes to act as opinion leaders, potentially amplifying or mitigating network polarization. This result implies that sharing mechanisms on online social networks can introduce imbalances in the opinion dynamics process compared to offline socialization, possibly resulting in higher levels of network polarization. These findings underscore the potential for online social networks to drive change and shape our social interactions. Finally, we validated our model with a dataset of real cascade on the Twitter platform and proved that the model is able to predict very well the real opinion dynamics. With our model, we can better understand the dynamics of online social networks and leverage this knowledge to create more effective communication strategies for businesses, organizations, and individuals alike.
II Related Works
The modeling of opinion dynamics has garnered the interest of researchers for decades, with this interest further amplifying with the emergence of online social networks. There are several classes of opinion dynamics models, and the literature on the topic is extensive [17]. Here, we concentrate on the class of averaging models, which posit that users adjust their opinions based on those of the people around them to mitigate social discomfort. Among these models, the FJ model [10] is one of the most studied [18, 19, 20, 21, 22, 23, 24, 25, 26, 14]. Its inner workings are explained in Seciton III-C. Recently, research on the FJ model has taken two main directions. The first involves optimization problems based on FJ [27, 28, 29, 21, 18, 19], whose main goal is to make interventions into the social network to achieve specific desired outcomes in the final opinions. For example, some investigate how to modify the initial opinion of a set of nodes to maximise the opinion discord [27] or minimize the overall opinion [28]. The second direction explores the possible extensions of the FJ model to capture diverse observed features in the online social networks [30, 31, 32, 33, 34]. In particular, these papers incorporate both attractive and repulsive relationships between users [31, 33], even assuming that the influence of a node on another can vary according to the homophily principle [34]. Furthermore, Bhalla et al. in [30] analyzes the polarization effect of FJ on time-varying networks and Chitra et al. in [32] extended the FJ to reproduce filter bubbles. Differently from these works, we want to investigate the effect of the classical FJ model in a static network when altering the social debate, which is regulated by the posting flow of news on online social networks.
To model the posting flow, we mainly take hints from the very popular independent-cascade model in [35] (see Section III-B for a description of how it works), which is able to reproduce realistic diffusion dynamics [36]. This seminal work has inspired many following studies and extentions [37, 38, 39, 40], which extend the classical model to include collaborative and/or antagonistic relationships among users or dynamic networks. The main goal of information diffusion is to study the rate of diffusion of viral contents depending on the parameters. Instead, we are not interested in the popularity of contents, but rather in their diffusion process. Hence, we used the original independent-cascade model and analyzed its impact on the opinion formation process. Similarly to opinion dynamics, the literature on information diffusion is extensive, so we refer the interested reader to dedicated surveys for more details [41].
The work closer to ours is [42] which proposes a model to estimate the inter-dependencies of viral contents and opinion formation. This model operates with two mechanisms during each time round.The first mechanism involves opinion updating, modeled using a relaxed version of the FJ model, which cannot reproduce polarization (for further details, see [14]). The second one is the proper interplay between information spreading and opinion adjustments. In this mechanism, the content spreads and its resulting popularity acts as a probability for nodes to change their innate opinion. In this model, the information cascade is not linked to people’s opinions when resharing the contents, nor does the flow of content sharing affect opinion evolution. In contrast, we assume that sharing content reflects peoples’ opinions and that the sharing process itself impacts opinion updating. Our goal is to understand how the propagation flow alters the FJ and investigate the impact on polarization. As far as we know, our work is the first approach to investigate this problem.
III Preliminaries
In this section, we will summarize the theoretical notions and notation we will use throughout the rest of the paper.
III-A The social network
We can describe an online social network platform as a graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$, where the $v$ nodes $\mathcal{V}$ represent users and the edges $\mathcal{E}$ represent social connections between them. In the most general form, the graph $G$ is directed and the edges $(i,j)$ represent a follow relationships. Specifically, $(i,j)\in\mathcal{E}$ means that $i$ follows $j$, thus $j$ influences $i$ as its activity appears in $i$’s news feed. The set of nodes that has an influence on $i$ is thus composed of $i$’s out-neighbours, which we will denote with $\mathcal{N}^{+}(i)$. Instead, the set of nodes that is influenced by node $j$ is the in-neighbourhood of $j$, which we indicate with $\mathcal{N}^{-}(i)$ [^1] Please note that, generally, the neighbourhood of a node corresponds to our out-neighbourhood.. It is worth noting that in the special case of an undirected node, $\mathcal{N}^{+}(i)$ and $\mathcal{N}^{-}(i)$ are the same. This social connection can arise, for example, when $i$ follows $j$ on platforms like X or Facebook or when they are friends on Facebook. The strength of this relationship is mathematically represented by the quantity $\hat{w}_{ij}$, which is the $(i,j)$-th entry of the $n\times n$ matrix $\hat{W}$ that we call social matrix. The weight that node $i$ gives to its out-neighbours represents the degree of influence of their opinions. It is provided by the $(i,j)$-th entry of the $n\times n$ matrix $W$, named the influence matrix $W$, which is row-stochastic. On the $i$-th row, for example, we can see how the influence on $i$’s opinion is distributed among all nodes in $i$’s network. In situations where additional data is lacking, the matrix $W$ can be derived from $\hat{W}$ through row-normalization, expressed as $w_{ij}=\frac{\hat{w}{ij}}{\sum_{j}\hat{w}{ij}}$. This process capitalizes on the insight that stronger relationships should carry greater weight in opinion formation. It’s worth noting that, with this mechanism, even in cases where the social graph exhibits symmetry, the influence matrix results asymmetric.
III-B Information cascade
Our model conceptualizes the act of posting social content following the independent cascade model [35]. In the independent cascade model, each node can be active or inactive. An initial set of active nodes $v$ have a single chance to activate their neighbours $w$, and each activation is successful with probability $p_{vw}$. The nodes $w$ that are successfully activated, become active at the next step and try to activate their neighbours as well. The process continues until no other nodes can be activated.
III-C Friedkin-Johnsen opinion dynamic model
The FJ opinion dynamic model is a mathematical model that describes how people’s opinions change over time due to their mutual influences. It does this by representing people’s opinions as real numbers and then modelling two opposing forces that affect how those opinions evolve. The first force is an initial bias that each person $i$ has, which is represented by $u_{i}$. The second force is a desire to reduce the discomfort people feel when their opinions differ from those around them, as described by the cognitive dissonance theory. Each person $i$ has a parameter $\lambda_{i}$, called susceptibilty, that determines how much weight they give to each of these forces. If $\lambda_{i}$ is 0, it means the person is completely stubborn and will never change its mind. If $\lambda_{i}$ is 1, it means the person is completely naive and will totally get over their initial bias. Node susceptibilities are generally represented by the diagonal matrix $\Lambda$, whose $i$-th diagonal element is equal to the susceptibility $\lambda_{i}$.
The FJ model is mathematically expressed as a simple update equation that is applied to each person’s opinion at discrete times. As discussed in our previous work [14], in the literature, the term “FJ model” refers to many variants of the original model. The update expression for the most general variant of FJ model is given by the following equation:
$x_{i}(t+1)=\lambda_{i}\sum_{j=0}^{v}w_{ij}x_{j}(t)+(1-\lambda_{i})u_{i}\qquad\forall i,$ (1)
where $w_{ij}$ is the $(i,j)$-th entry of matrix $W$. The first term of the sum represents the movement of $i$’s opinion towards the opinion of its neighbours, weighted by their influence degree. The second term represents instead the anchoring to the initial prejudice. Under general assumptions, the model converges, as summarized in the following theorem (proved in [23]).
Theorem 1.
[FJ convergence] If the graph is strongly connected and $\Lambda\neq I$ (i.e., at least one agent partially anchored to its initial opinion exists) [^2] Please note that when $\Lambda=I$, the FJ model reduces to the French-DeGroot model [11], which is a simpler model that has been separately studied in the literature. then the FJ model is convergent, and it converges to the stable opinion vector given by:
$\bm{z}=(I-\Lambda W)^{-1}(I-\Lambda)\bm{u},$ (2)
where $I$ is the identity matrix.
This version of the FJ model is synchronous, meaning that all nodes update their opinion simultaneously. However, in reality, people engage in asynchronous interactions, so opinions do not evolve simultaneously. In other words, we do not engage with all our social contacts all at once. To capture this, the FJ model also has an asynchronous version [43] that, as we will discuss later, has been proven to converge to the synchronous FJ model of Equation (1). In the asynchronous model, an edge $(i,j)$ is selected uniformly at random at each time step and the opinion of only one of the two nodes (say $i$, selected at random) is updated, according to the following equation
$x_{i}(t+1)=h_{i}\left[(1-\gamma_{ij})x_{i}(t)+\gamma_{ij}x_{j}(t)\right]+(1-h_{i})u_{i},$ (3)
while the rest of the nodes maintain their opinion. Here matrices $H=\{h_{ij}\}$ and $\Gamma=\{\gamma_{ij}\}$ replace $\Lambda$ and $W$ of the synchronous model, and are different from them. In fact, to guarantee the convergence of the model in Equation (3) to that in Equation (1), the formulation has to take into account the probability of node $i$ being randomly selected, which, in turn, depends on the degree of nodes. Since high-degree nodes are involved in more edges, when selecting edges at random high-degree nodes have a higher probability of being selected. For this reason, $H$ and $\Gamma$ should be defined from $\Lambda$ and $W$ as follows:
$h_{i}=1-\frac{1-\lambda_{i}}{d_{i}}$ (4)
$\gamma_{ij}=\left\{\begin{array}[]{ll}\frac{d_{i}(1-h_{i})+h_{i}-(1-\lambda_{i}w_{ii})}{h_{i}}&\textrm{if }i=j\\ \frac{\lambda_{i}w_{i}j}{h_{i}}&\textrm{if }i\neq j\end{array}\right.$ (7)
where $d_{i}$ is the degree of node $i$. In [43], Frasca et al. establishes the ergodic convergence of the asynchronous FJ algorithm defined by equations (3) to the classical FJ. Notably, while the individual state variables generally do not converge in the pointwise sense and often exhibit persistent oscillation, their time-averaged values converge asymptotically to a stable limit, which, in our case, coincides with the stable point of the FJ.
III-D Polarization
In our previous work [14], we reviewed and classified the most popular polarization metrics used in the related literature. These metrics, given an opinion vector, measure polarization according to different semantics. They can be summarized in the following definition.
Definition 1.
For an opinion $\bm{x}=(x_{i})\in[-1,1]^{v}$ the following polarization indices are defined:
$NDI(\bm{x})$ $=\sum_{(i,j)\in\mathcal{E}}w_{ij}(x_{i}-x_{j})^{2}$ (8)
$GDI(\bm{x})$ $=\sum_{i,j\in\mathcal{V}:i<j}(x_{i}-x_{j})^{2}$ (9)
$P_{1}(\bm{x})$ $=\sum_{i\in\mathcal{V}}(x_{i}-\bar{\bm{x}})^{2}=\|\bm{x}-\bar{\bm{x}}\|_{2}^{2}$ (10)
$P_{2}(\bm{x})$ $=\frac{1}{|\mathcal{V}|}\sum_{i\in\mathcal{V}}x_{i}^{2}=\frac{1}{|\mathcal{V}|}\|\bm{x}\|_{2}^{2}$ (11)
$P_{3}(\bm{x})$ $=\sum_{i\in\mathcal{V}}x_{i}^{2}=\|\bm{x}\|_{2}^{2}$ (12)
$P_{4}(\bm{x})$ $=\sum_{i\in\mathcal{V}}|x_{i}|=\|\bm{x}\|_{1}$ (13)
In the context of measuring polarization in social networks, the metrics $NDI$, $GDI$, and $P_{1}$ quantify polarization based on the degree of disagreement among individuals’ opinions. On the other hand, metrics $P_{2}$, $P_{3}$, and $P_{4}$ consider polarization to occur when opinions are significantly distant from a neutral opinion, which is assigned a value of 0.
A dynamic opinion model is considered polarizing with respect to a given metric $\Phi$ if there exists at least one initial opinion vector $\bm{u}$ such that the corresponding final opinion $\bm{z}$ exhibits greater polarization, i.e., $\Phi(\bm{z})>\Phi(\bm{u})$. The quantity
$\Delta_{\Phi}(\bm{u})=\Phi(\bm{z})-\Phi(\bm{u})$ (14)
is referred to as the polarization value and, by definition, is positive if and only if the model is polarizing with respect to the metric $\Phi$. Our previous research has demonstrated that the FJ model is in fact polarizing, and a formula for obtaining polarizing initial opinion vectors for the metrics $P_{2}$, $P_{3}$, and $P_{4}$ have been identified. For what concerns vectors that maximise $P_{2},P_{3}$ polarization, they are correlated to the spectral properties of matrix $H=(I-\Lambda W)^{-1}(I-\Lambda)$ which transforms the initial opinions $\bm{u}$ into the final opinions $\bm{z}$ as described in Equation (2). In particular, we identified $\bm{u}_{B_{2}(1)}$ maximises polarization on the $L_{2}$-ball of radius 1 and $\bm{u}_{B_{2}(t)}=t\cdot\bm{u}_{B_{2}(1)}$ is its scalar multiple that maximises polarization on the $L_{2}$-ball of radius $t$. Furthermore, $\bm{u}_{V_{>1}}$ maximises polarization on the subspace of the specific $H$’s eigenvectors and $\bm{u}_{\max_{P_{2,3}}}$ on the entire domain. For what concerns vectors that maximise $P_{4}$ polarization, we identified $\bm{u}_{B_{1}(1)}$ that maximises polarization on the $L_{1}$-ball of radius 1 and $\bm{u}_{\max_{P_{4}}}$ on the entire domain. Unfortunately, while vectors $\bm{u}_{B_{2}(1)}$, $\bm{u}_{B_{2}(t)}$ and $\bm{u}_{B_{1}(1)}$ can be always found, $\bm{u}_{V_{>1}}$, $\bm{u}_{\max_{P_{2,3}}}$ and $\bm{u}_{\max_{P_{4}}}$ are often impossible to be computed since they are the solutions of NP-hard problems. However, $\bm{u}_{V_{>1}}$ can often be estimated with a heuristic. Please refer to [14] for further details. In Section V we will examine the impact of information cascades when the opinion dynamic process starts with users featuring these polarizing opinions. Notably, we will observe that FJC model induces significantly greater polarization than the FJ for most of the scenarios studied.
IV The Friedkin-Johnsen on Cascades model
In this section, we will discuss the FJC (Friedkin-Johnsen on Cascades) model. The FJ model is an opinion dynamic model that explains how people form their opinions in the offline world, where face-to-face interactions significantly influence the depth of relationships and consequently the shared understanding of opinions. This is the unique mechanism that triggers the tension resolving process, which is the brain’s reaction to cognitive dissonance. Information cascades, on the other hand, describe how information is spread through online social networks, where interactions are encoded as engagement with each other’s posts. They characterise how a single post by a source can go viral and inundate the entire network. Despite belonging to two different and separate research areas, opinion dynamics and information cascades are intricately linked. In online social networks, opinions are expressed in a cascade manner, by subsequent exposure to other contents/opinions. It is reasonable to expect that this unique exposure to others’ posts influences how we update our opinions. The amplification of virality in social networks can disrupt the balance of real-life discussions in offline networks, thus strengthening the role of central nodes. This is a crucial insight that has far-reaching implications for understanding opinion formation and information dissemination in online social networks.
IV-A Revisiting the independent cascade model
The standard independent cascade model needs to be slightly adjusted to integrate it into the FJC framework. To this aim, we identify three new node states, which replace the previous ones: unaware, aware and spreader. A node is unaware when it has not been exposed to the post yet. As soon as it “sees” the post from a neighbour, the node becomes aware and then it updates its opinion as described in the following sections. With a probability $\theta_{s,i}$ it also reshares the post, and, when it does, the node moves into the spreader state. An illustration of the possible transition of states for a single node is provided in Figure 1. We assume (A1) that the content being shared does not express an opinion on its own, but rather the opinion is conveyed only by the person sharing it along with a comment. This assumption allows us to avoid considering the post as an external factor that can modify opinions and instead to focus on studying how the simple cascade dynamic affects opinion formation in comparison to the FJ scenario.
Figure 1: States, transitions and actions of the FJC model for a single node. The blue boxes indicate the states and the orange arrows with corresponding probabilities indicate the transitions between states. The purple arrows and boxes indicate the actions made within the corresponding states.
A post may become obsolete and lose its relevance over time. When this occurs, nodes cease to propagate it, even if some nodes have not been reached. We capture this effect by defining a threshold parameter $L$, known as the interest threshold. Specifically, $L$ represents the maximum hop distance that a post can travel from its source to its destinations. Thus, once a post has reached $L$ hops, its propagation ceases. It is important to note that if this threshold is greater than or equal to the diameter of the network (defined as the longest shortest path between any two nodes in the graph $\mathcal{G}$), all posts will always propagate throughout the entire network, rendering the parameter $L$ ineffective in influencing the resulting cascade.
In our model, an individual identified as the source disseminates a piece of news to the online social network. In the rest of this paper, we streamline our discussion by initially constructing the model under the assumption that each post originates from a single seed node. However, it is worth noting that the model’s flexibility allows for straightforward extension to accommodate multiple sources, akin to the original information cascade model. In fact, including multiple sources simply affects the first level of the information cascade tree, while the rest remains unchanged. We assume (A2) that once the post is shared the first time by the seed node $r$, it spreads outwards like a tree, from the original source to the first, second, third, and subsequent in-neighbours, which we will refer to as the nodes in the first, second, third, … layer. This assumption means that the opinions of parent nodes represent the first exposure a node has in its news feed, thereby solidifying their influence. Following this initial encounter, we maintain that nodes typically become disinterested in posts shared by others. Thus, we will refer to the $l$-th in-neighbourhood as the $l$-th layer or layer $l$ and we will indicate it with $\mathcal{O}_{l}^{(r)}$. Keeping in mind that $L$ is the interest threshold, the entire tree that comprehends all nodes in the aware state of the cascade is denoted as the sequence $\mathcal{O}^{(r)}=\left(\mathcal{O}_{l}^{(r)}\right)_{l=0}^{L}$ of all layers.
A post $p$ generated at node $r$ spreads into the network over the tree $\mathcal{O}^{(r)}$. However, not all nodes in this tree receive the content, because there is a probability $\theta$ that each node will not reshare the post. The subtree of $\mathcal{O}^{(r)}$ that includes only nodes that have received the post is denoted as $\mathcal{T}^{(r,\theta)}=(\mathcal{T}^{(r,\theta)}_{l})_{l=0,\dots,L}$, and it holds that $\mathcal{T}^{(r,\theta)}_{l}\subseteq\mathcal{O}^{(r)}_{l}$ for each $l$. $\mathcal{T}^{(r,\theta)}$ is a stochastic realization of the tree $\mathcal{O}^{(r)}$, as indicated by the superscript $\theta$. In Figure 2, there is a simple example of a network and a corresponding propagation tree.
(a) Network
(b) $\mathcal{T}^{(1|\theta)}$
Figure 2: A toy example of a network in (a) and the corresponding propagation tree for $r=1$ in (b). Nodes $2$ and $3$ update their opinion deterministically, while the others, do it with probability $\theta$.
We can describe it in formal terms in the following way. We refer to the cascade generated by the post $p$ originating in node $r$ as the collection of nodes in the tree $\mathcal{T}^{(r,\theta)}$ and we will denote it with $\mathcal{C}^{(r)}$. Please observe that nodes in layer 1 always belong to the cascade because they receive the post directly from the source, while those of layer $l\geq 2$ belong to the cascade if and only if they receive the post from nodes in layer $l-1$.
Let us indicate with $S_{i}$ the random variable modelling the event of resharing a post for a node $i\neq r$ (it takes the value 1 when the post is reshared, 0 otherwise). Then, the tree $\mathcal{T}^{(r,\theta)}$ is described as:
$\left\{\begin{array}[]{l}\mathcal{T}^{(r,\theta)}_{0}=\{r\}\\ \mathcal{T}^{(r,\theta)}_{1}=\mathcal{N}^{-}(r)\\ \mathcal{T}^{(r,\theta)}_{l+1}=\left(\bigcup\limits_{\{i\in\mathcal{T}^{(r,\theta)}_{l}:S_{i}=1\}}\mathcal{N}^{-}(i)\right)\setminus\mathcal{T}^{(r,\theta))}_{l}.\end{array}\right.$ (15)
Please, remember that $\mathcal{N}^{-}(i)$ refers to the in-neighbours of node $i$.
For our convenience, for each node $i\in\mathcal{O}^{(r)}_{l}$ we denote with $\varphi(i)$ the set of nodes who actually influence $i$’s opinion in the stochastic realization of the cascade, i.e the nodes in layer $l-1$ that are in $i$’s out-neighbour and reshare the post. In other terms:
$\varphi(i)=\left\{j\in\mathcal{N}^{-}(i):j\in\mathcal{T}^{(r,\theta)}_{l-1}\wedge S_{j}=1\right\}$ (16)
We will refer to this set as the predecessors of $i$.
IV-B The cascade-driven Friedkin-Johnsen
We will now discuss how to integrate the independent cascade model described in Section IV-A with the FJ model from Section III-C, resulting in a FJ variant (FJC) where opinions are influenced by the sharing of posts. In this model, opinions are influenced by the sharing of posts among nodes. The FJC model operates iteratively, following specific states and transitions for each node (see Figure 1). At the start of each cascade, every node, except for the designated seed node $r$, begins in the unaware state. During each iteration, nodes $i$ that are exposed to a post from another node $s$ undergo two consecutive actions. First, these exposed nodes transition to the ”aware” state and update their opinions regarding the post. Following this, with a probability $\theta_{s,i}$, these nodes enter the spreader state, sharing the post again if they belong to a layer $l$ that is smaller than the maximum layer $L$. If a node does not meet this condition, the propagation of the post stops, and a new cascade is initiated. We will examine the spreading process of $N$ posts, denoted with $p_{1},\dots,p_{N}$. We make an assumption (A3) that posts propagate one at a time, meaning that simultaneous spread of two posts is not possible. While this assumption is stringent, it aids in designing the model .In future work, we will explore ways to relax this assumption and understand its impact on the model. However, in Section V-C, we demonstrate that the model is robust in real-world settings.
Now, we want to discuss the details of the opinion updating operation. Please observe that, since we want to update opinion layer by layer, we need an asynchronous version of the FJ, similar to that of Equation (3).
We focus on the spreading process of the $n$-th post $p_{n}$ generated at node $r_{n}$ which involves nodes in the cascade $\mathcal{C}^{(r_{n})}$. As the cascade starts, the opinions of nodes are equal to $\bm{u}=\bm{x}(0)$ if $n=1$, i.e. if we are considering the spreading of the first post. Otherwise, they are equal to the final opinion of nodes as the previous post $p_{n-1}$ completed spreading. Every time a node $j$ in the cascade reshares the post $p_{n}$ at a time $t$, it conveys its current opinion $x_{j}(t)$ together with the reposting. A node $i$ in layer $l$ can receive the post (which, as explained above, is a proxy for a node’s opinion) from one or more predecessors $j$. This triggers the update of node $i$’s opinion, which is carried out according to the following update rule:
$x_{i}$ $(t)=$
$=\frac{1}{|\varphi(i)|}\Bigg{\{}\sum_{j\in\varphi(i)}h_{i}\left[(1-\gamma_{ij})x_{i}(t-1)+\right.$
$\qquad\left.+\gamma_{ij}x_{j}(t-1)\right]+(1-h_{i})u_{i}\Bigg{\}},$ (17)
where $h_{i}$ and $\gamma_{i}$ are defined as in Equations 4 and 5, and $\varphi(i)$ is defined in the Equation (16). The above update rule factors in the average total influence on node $i$ by all its in-neighbours that have reshared the post, hence shared their opinion with $i$. Note that, differently from the asynchrounous FJ update in Equation (3), node $i$ assesses the opinions of its predecessors all at once, not in a pairwise fashion.
Differently from the FJ model in Equation (1), we update opinions node by node, following the order provided by the layer of tree $\mathcal{T}^{(r_{n},\theta)}$. Once the opinions of a single layer are updated, the post spreads again. In Algorithm 1 we summarize the entire FJC process.
Algorithm 1 FJC
1:$\bm{R}$ (vector of post seeds), $\bm{u}$ (initial opinion vector)
2:$\bm{x}$ (final opinion vector)
3:$\bm{x}\leftarrow\bm{u}$triangleright initialize the opinion vector
4:for $r$ in $\bm{R}$ do
5:     for $l$ in $\{1,\dots,L\}$ do $\triangleright$ iterate over layers
6:         for $i$ in $\mathcal{T}^{(r,\theta)}_{l}$ do
7:              $x_{i}\leftarrow$ Equation  (17) $\triangleright$ update opinions
8:         end for
9:     end for
10:end for
V Performance Evaluation
In this section, we analyze the performance and validate the FJC model across different social networks. To this end, we conducted two distinct sets of experiments. The first set (Sec. V-B) aimed to compare the opinion dynamics produced by the FJC model with those of the classical FJ model. For this purpose, we used three social networks—two real-world and one synthetic—and generated synthetic cascades on top of them. Specifically, we selected 15 seed nodes per graph to initiate the simulated cascades. The second set of experiments (Sec. V-C) focused on validating the FJC model using real-world cascade data. Here, we evaluated how well the model can approximate opinion evolution based on traces from an online social network, where assumptions (A2) and (A3) do not always hold. Specifically, (A2) posits that cascades spread in a tree-like fashion from the seed node, layer by layer, and (A3) assumes that only one post propagates at a time. Assumption (A1), however, could not be relaxed, as it underpins the model’s formulation of opinion dynamics — it assumes that the content itself is opinion-neutral, and opinions are conveyed solely by the user sharing the post. In the following, we provide a detailed description of the experimental settings and discuss the results obtained.
V-A Common experimental settings
All networks used are unweighted, so the social matrix $\hat{W}$ corresponds to the adjacency matrix. As described in Section III-A, the influence matrix is derived by row-normalizing $\hat{W}$: $w_{ij}=\frac{\hat{w}{ij}}{\sum_{k}\hat{w}{ik}}$. Following our previous work [14], we considered three strategies for setting node susceptibility. The first two rely on PageRank centrality, a measure that effectively captures node influence [23]. For each node $i$ with PageRank score $\mathrm{P}_{i}$, we assigned $\lambda_{i}$ either the rescaled value of $\mathrm{P}_{i}$ or $\mathrm{P}_{i}^{-1}$, both mapped to the interval $(0,1)$, thus giving higher susceptibility to more (or less) central nodes. In the third strategy, all nodes were assigned a fixed susceptibility value of 0.6. These variations allow us to explore how susceptibility interacts with node centrality. Finally, since our primary goal is to study the impact of cascades on polarization, we initialized opinions using the polarizing vectors defined in Section III-D, ensuring a consistent and controlled setup for measuring polarization dynamics.
V-B FJC vs FJ – with synthetic cascades
V-B1 Dataset description and experimental setting
In this section, we evaluate the performance of the FJC model through simulations on three social networks: the Karate Club graph [44], a Facebook graph [45], and a synthetic Barabási-Albert network [46]. The Karate Club graph is a well-known small network with 34 nodes, while the Facebook graph includes 1519 nodes from the largest connected component of a 4039-node snapshot. The Barabási-Albert network, which captures the scale-free structure of real social networks, was generated with 100 nodes and parameter $k=6$, meaning each new node connects to 6 existing nodes. Figure 3 shows visualizations of all three graphs.
Figure 3: Networks used for the performance evaluation. In (a) and (b) are two real social networks; in (c) a synthetic network.
Since cascades are simulated, we randomly selected 15 seed nodes per network to initiate the posts. We set the resharing probability $\theta_{s,i}$ uniformly to a single value $\theta$ for all nodes, testing three scenarios: low diffusion ($\theta=0.01$), moderate diffusion ($\theta=0.8$), and maximum diffusion ($\theta=1$), where nodes always reshare posts. For each combination of network, susceptibility setting, reposting probability $\theta$, and initial opinion values, we performed 1000 simulation runs.
V-B2 Opinion evolution under FJC vs FJ
We begin by analyzing how cascades influence final opinions under the FJC model compared to the classical FJ model. In Figure 4, each arrow represents a node’s opinion shift—from its initial opinion (prejudice) to its final FJC opinion. The black dot indicates the corresponding final opinion under the FJ model. Panels A and B show cases where susceptibility is proportional or inversely proportional to node centrality, respectively; panel C shows uniform susceptibility ($\lambda=0.6$). Due to space constraints, we report results for $\theta=0.01$ and $\theta=0.8$ (shown in subpanels (i) and (ii)), and for the polarizing vectors $\bm{u}_{B_{1}(1)}$, $\bm{u}_{B_{2}(1)}$, $\bm{u}_{B_{2}(t)}$ and $\bm{u}_{V_{>1}}$, and $\bm{u}{V_{>1}}$. The full results are available in the SI Appendix.
We start with panels C of Figure 4, where all nodes have equal susceptibility. Compared to the FJ model, highly central nodes in the FJC model tend to retain their initial opinions more strongly—visible from the shorter arrows relative to the distance to the FJ outcomes (black dots)—especially for $\theta=0.01$ (Ci) and $\theta=0.8$ (Cii). This leads to a wider spread of final opinions among central nodes. In contrast, less central nodes are more influenced and their opinions shift closer to those of central ones. Overall, cascades amplify the influence of network topology: central nodes act as opinion anchors, while peripheral nodes become more susceptible. The reposting probability $\theta$ modulates this effect—low values ($\theta=0.01$) limit opinion spread, while higher values ($\theta=0.8$) increase it. When susceptibility is inversely proportional to centrality (panels B), central nodes are already stubborn by design in both models, leading to similar outcomes. Nonetheless, cascades still shift the opinions of less central nodes more in the FJC model, as they become increasingly influenced by central ones. Finally, with susceptibility proportional to centrality (panels A), network topology and susceptibility act in opposing directions. Peripheral nodes are nearly immobile due to low susceptibility, while central nodes—despite being more susceptible—change little due to their structural position. As a result, final opinions under FJC remain close to initial ones, more so than in the FJ model. This highlights how, in cascade-driven dynamics, centrality outweighs susceptibility, with central nodes maintaining influence regardless of the surrounding opinion mass.
Figure 4: Polarizing opinions vs. node centrality. Each arrow represents a node, starting at its initial opinion (prejudice) and ending at its final opinion under the FJC model. Arrow color indicates node susceptibility; black dots show the final opinions under the FJ model. (A) $\lambda_{i}\propto\mathrm{P}_{i}$; (B) $\lambda_{i}\propto\mathrm{P}_{i}^{-1}$; (C) $\lambda_{i}=0.6$; (i) $\theta=0.01$; (ii) $\theta=0.8$. In each panel, left plots correspond to $P_{4}$-polarization vectors, right plots to $P_{2}$/$P_{3}$ vectors. Dotted and dashed lines show the average initial and final FJC opinions.
V-B3 Polarization
Figure 5 shows how opinion dynamics under FJC affect polarization compared to the FJ model, measured as the change in polarization (Equation (14)) between final and initial opinions. We report results for each susceptibility setting—proportional (A), inversely proportional (B), and uniform (C)—and for each $\theta$ value (differently colored bars for FJC). Due to space constraints, we show only the polarization metric maximized by each polarizing vector in the FJ model: $P_{4}$ for $\bm{u}_{B_{1}(1)}$ and $P_{2}$ for the others. Full results are in the SI. We first consider FJC with $\theta=0.8$ and $\theta=1$ (blue and purple bars). When susceptibility is uniform (panel C), FJC consistently yields higher polarization than FJ, as central nodes remain influential and pull others toward their opinions. A similar pattern, though less pronounced, appears with inverse-centrality susceptibility (panel B), since less central nodes become more vulnerable under FJC. Interestingly, when susceptibility is proportional to centrality (panel A), FJC often reduces polarization compared to FJ. Here, central nodes are nearly immovable, and peripheral nodes, unable to influence them, shift closer to reduce cognitive dissonance—flattening the opinion landscape. An exception occurs in the Facebook network with $\bm{u}_{B_{1}(1)}$, where the initial opinion aligns with a central node that remains firm, attracting others and increasing polarization. For $\theta=0.01$ (green FJC bars), the cascade effect is weak due to limited reposting, resulting in polarization trends similar to those at higher $\theta$ but with lower magnitude. In some cases, the effect even reverses, leading to depolarization—primarily driven by the few nodes directly connected to the sources, which are the only ones likely to update their opinions. For instance, in the Karate network with proportional susceptibility and the vector $\bm{u}_{B_{1}(1)}$, most neighbors of the sources have low susceptibility and thus remain unchanged. The only significant shift comes from the source node itself, which lowers its opinion from 1, reducing overall polarization as measured by $P_{4}$.
Figure 5: Polarization of FJ and FJC (with different $\theta$ values) for (A) $\lambda_{i}\sim\mathrm{P}_{i}$, (B) $\lambda_{i}\sim\mathrm{P}_{i}^{-1}$, (C) $\lambda_{i}=0.6$.
V-C FJC with real cascades
V-C1 Dataset description and experimental setting
We validate the model using a real-world dataset from Bhowmick et al.  [47], which contains Twitter (now X) data related to the 2015 Nepal earthquake. This dataset is well-suited for our analysis as it includes both the follower network and tweet/retweet histories. Unlike the synthetic networks in Sec. V-B, this is a directed graph, with multiple connected components. Since FJ-based models operate within a connected structure, we extract the largest strongly connected component (LSCC). Some cascades span disconnected components, suggesting hidden links, so we restrict our analysis to cascades occurring entirely within the LSCC. The resulting graph contains 220,974 nodes, 16,772, 693 edges, and 32,313 cascades. Initial opinions are set using the polarizing vectors described in Section III-D. However, due to the large size of the network, even optimized methods could not compute these vectors over the full LSCC. We therefore define two subgraphs:
  • Nepal cut degree: includes only nodes with degree in the range $[150,400]$ (where the node degree spans from 2 to 53718 in the original LSCC), balancing network activity and computational feasibility. We extract the LSCC and retain only the cascades involving its nodes.
  • Nepal top cascades: constructed from nodes involved in the 500 longest cascades using a heuristic for the directed Steiner Tree problem [48] to ensure that all the nodes involved in these cascades are in the subgraph.
Table I summarizes the two subgraphs. The Nepal cut degree graph has more nodes but fewer edges, while Nepal top cascades is denser and includes more cascades (hence making it more likely that assumption (A3) is violated).
TABLE I: Summary of the two real-world Nepal datasets subgraphs used for validation
#nodes #edges #cascades max cascade length
Nepal cut degree 30662 737541 4111 4
Nepal top cascades 26321 4222064 6083 21
Each dataset $D$ provides a sequence of tuples $\mathcal{C}_{D}=\big{(}(t_{1},p_{1},i_{1}),(t_{2},p_{2},i_{2}),\dots\big{)}$, where node $i_{k}$ shares post $p_{k}$ at time $t_{k}$, representing the temporal evolution of cascades. We assume that this sequence of reshares drives the opinion dynamics in the real world: each reshare is treated as an expression of opinion. At time $t_{k}$, when node $i_{k}$ shares a post, its followers update their opinions according to Equation (17). This process relies solely on assumption (A1); assumptions (A2) and (A3) are not used. We refer to this approach as real opinion dynamics. It is a data-driven heuristic and cannot be analyzed analytically.
Our goal is to evaluate how well this real dynamic is captured by the proposed FJC model (we also include the classical FJ model for reference). For the FJC model, we need to estimate the resharing probability $\theta_{i}$ for each node. Inspired by weighted cascade models [35, 49], we define $\theta_{i}$ as the ratio between the number of posts reshared by node $i$ and the number of posts it has seen from its followees. Table II shows the distribution of $\theta_{i}$, which is generally low, indicating that most nodes rarely reshare posts. Unlike the real dynamics, the FJC model processes cascades in a tree manner and sequentially. Specifically, posts are propagated from the sources to the network layer-by-layer, in accordance with assumption (A1) and each cascade starts only after the previous one ends, in accordance with assumption (A3).
TABLE II: Statistics of followers, followed and $\theta$ in both Nepal subgraphs
Nepal cut degree
Min. 1st Qu. Median Mean 3rd Qu. Max. N.A.
Followers 1.00 8.00 19.00 24.05 34.00 182.00 -
Followees 1.00 13.00 21.00 24.05 32.00 135.00 -
$\theta$ 0.000 0.000 0.000 0.001 0.000 1.000 8807
Nepal top cascades
Min. 1st Qu. Median Mean 3rd Qu. Max. N.A.
Followers 1.0 23.0 59.0 160.4 148.0 11022.0 -
Followees 1.0 53.0 109.0 160.4 204.0 6851.0 -
$\theta$ 0.0000 0.0000 0.0000 0.0156 0.0000 1.0000 681
V-C2 Opinion evolution of real opinion dynamics vs FJC and FJ
Due to the large size of the networks, we do not visualize the polarizing vectors as in Figure 4, since the plots would be unreadable. Instead, our main objective with these datasets is to evaluate the robustness of the FJC model against the real opinion dynamics. We do so by comparing the distributions of opinion shifts—i.e., the differences between final and initial opinions—for the real evolution (labeled Real) and the predictions of FJC and FJ. As shown in Figure 6, FJC closely matches the real dynamics in the Nepal cut degree dataset (the two curves are nearly indistinguishable) and shows reasonable agreement in the Nepal top cascades dataset. This is expected, as longer cascades are more likely to violate assumption (A3), reducing FJC’s accuracy. Nevertheless, FJC remains surprisingly robust even in this more challenging scenario.
Figure 6: ECDF of the opinion shift of nodes, i.e. final opinion minus initial opinion, in the three different opinion dynamics. (A) $\lambda_{i}\sim\mathrm{P}_{i}$, (B) $\lambda_{i}\sim\mathrm{P}_{i}^{-1}$, (C) $\lambda_{i}=0.6$.
V-C3 Polarization
Figure 7 shows polarization outcomes across different settings. Both cascade-based dynamics—FJC and Real—exhibit very low or even negative polarization, in sharp contrast to the FJ model. This suggests that, in this case, cascades tend to reduce polarization, seemingly contradicting the findings from synthetic networks in Section V-B. However, this can be explained by the fact that, as shown in Table II, most nodes have $\theta$ values close to zero, indicating minimal reposting activity. These results are therefore consistent with our earlier observation: when $\theta$ is small, opinion dynamics are largely inactive, preventing polarization from developing.
Figure 7: Polarization value of the different vectors in the corresponding metrics in the three different opinion dynamics. (A) $\lambda_{i}\sim\mathrm{P}_{i}$, (B) $\lambda_{i}\sim\mathrm{P}_{i}^{-1}$, (C) $\lambda_{i}=0.6$.
VI Conclusions
This study explores the dynamics of opinion formation and dissemination within online social networks, particularly under the influence of information cascades. Specifically, the paper proposes the Friedkin Johnsen on Cascade model (FJC), which integrates opinion dynamics and information cascades. To the best of our knowledge, this is the first model that allows for the investigation of how content dissemination can affect opinion formation. We evaluate the model using both synthetic and real social networks and highlight that the cascades reinforce the strength of central nodes in influencing opinions. These central nodes tend to be less vulnerable to opposing opinions, making it unlikely for them to change their views even when pressured by a large majority. This occurs when content is free to circulate within the network. Conversely, when participation is low, content cannot go viral, disrupting the opinion dynamics. The model builds upon certain assumptions regarding content dissemination and opinion evolution. To validate the FJC model, we utilized a real cascade dataset. Our results demonstrate that the FJC model is robust and can approximate final opinions accurately, even when many of the underlying assumptions of the model are violated.

### v2

@@
 As OSNs become increasingly prominent sources for news dissemination, individuals often encounter content that influences their opinions through both direct interactions and broader network dynamics.
-In this paper, we propose the Friedkin-Johnsen on Cascade (FJC) model, which is, to the best of our knowledge, is the first attempt to integrate information cascades and opinion dynamics, specifically using the very popular Friedkin-Johnsen model.
+In this paper, we propose the Friedkin-Johnsen on Cascade (FJC) model, which, to the best of our knowledge, is the first attempt to integrate information cascades and opinion dynamics, specifically using the very popular Friedkin-Johnsen model.
 Our model, validated over real social cascades, highlights how the convergence of socialization and sharing news on these platforms can disrupt opinion evolution dynamics typically observed in offline settings.
@@
 I Introduction
-Socialization represents a fundamental aspect of human behavior, with empirical evidence consistently demonstrating the correlation between human health and social engagements [1, 2, 3], underscoring how isolation has serious consequences on both physical and mental health conditions [4, 5, 6], and could potentially lead to premature mortality [7].
-The advent of online social networks (OSNs) has significantly reshaped how individuals fulfill their social needs, offering unprecedented opportunities to connect with others.
+Online social networks (OSNs) have fundamentally transformed how individuals communicate, share information, and form opinions in the digital age.
+Serving both as platforms for social interaction and as primary sources of real-time news, OSNs facilitate unprecedented levels of connectivity and information dissemination [1].
 Additionally, OSNs have emerged as prominent sources of real-time news dissemination, with a substantial proportion of the population relying on these platforms for information [1].
@@
 These dynamics underscore the complex interplay between socialization, news consumption, and information dissemination in the digital era, warranting further empirical investigation and theoretical exploration within interdisciplinary frameworks.
-The dynamics of opinion formation represent a multifaceted research area situated at the intersection of various disciplines, including sociology, economics, psychology, control theory, computational sociology, and physics.
+Understanding the mechanisms behind these phenomena requires an integrated study of how opinions evolve through social influence and how information cascades propagate within these networks.
+The dynamics of opinion formation represent a multifaceted research area situated at the intersection of various disciplines, including sociology, economics, psychology, control theory, computer science, computational social science, and physics.
 Numerous models have been proposed to describe the mechanisms underlying opinion dynamics, aiming to capture observed theoretical and empirical phenomena.
@@
 It has demonstrated the ability to replicate various observed mechanisms in social networks, including the phenomenon of wisdom of the crowd [6] and the emergence of polarization [7] and recently, it has also been proposed as the building block to model multilateral international negotiation processes [8].
-The FJ model has undergone rigorous validation in offline scenarios, where opinions evolve reflecting real-world interactions and debates on various topics.
-However, the dynamics change when socializing online, as we are exposed to friends’ opinions not through direct interaction but via their public activity on social media platforms, such as posting or commenting on news articles.
-Consequently, we may frequently encounter cognitive dissonance toward opinions that are expressed more frequently online.
-These online opinions often garner larger-scale attention and spread more intensely than their offline counterparts, potentially heightening the impact of social opinion exchange in digital social environments.
-These opinions are conveyed via textual posts, resharing of news items, or images.
-The dynamics of how content items are shared and spread online are studied under the name of information cascades [16].
+The Friedkin-Johnsen (FJ) model has been validated in offline settings, where opinion evolution reflects direct interpersonal interactions and deliberations across various topics [9, 10].
+However, online social environments introduce distinct mechanisms of opinion exchange: individuals are exposed to their contacts’ expressed opinions indirectly, primarily through public social media activities such as posts, comments, and shares [11, 12].
+This mode of exposure frequently involves repeated encounters with viewpoints from a broader and more heterogeneous audience, which can heighten cognitive dissonance by presenting conflicting information [13, 14].
+In online environments, opinion exposure is both asynchronous and large-scale, allowing conflicting signals to endure over time and reach diverse audiences.
+At the same time, empirical studies show that controversial or polarizing opinions propagate with exceptional speed and breadth compared to offline interactions, amplifying their influence on collective attitudes [15, 16].
+These opinion expressions propagate through various content formats, including textual posts, news resharing, and multimedia, whose spread dynamics are effectively modeled by the theory of information cascades [17].
 Opinion dynamics and information cascades operate at distinct granularity and temporal scales.
@@
 Conversely, information cascades emphasize the sharing mechanism of content without addressing the effects of such sharing.
-However, given that reposting online content received from friends is a common online interaction, it is reasonable to expect an intertwining of opinion dynamics and information cascades.
-Despite this, the two processes have thus far been treated separately in the literature.
+Importantly, reposting online content received from friends (e.g., retweets) is one of the most common forms of interaction in digital platforms.
+Empirical evidence shows that in a complete 24-hour snapshot of Twitter activity, only 20.8% of tweets were original, while 50.7% were retweets, underscoring that resharing dominates user behavior [18].
+This widespread resharing suggests a natural intertwining between opinion dynamics and information cascades: as users repeatedly encounter and propagate content, microscopic acts of resharing can influence macroscopic opinion evolution.
+Despite this evident overlap, the two processes have thus far been treated separately in the literature.
 In this work, we set out to fill this gap and establish an integrated model whereby the microscopic act of content sharing bears an effect in the macroscopic phenomenon of opinion evolution in online social networks.
-To address this issue, we propose a novel model that integrates the FJ model with information cascades.
-Specifically, we hypothesize that opinions on a selected topic are expressed following the dissemination flow of online posts.
-Initially shared by certain sources, the post propagates as a cascade to nodes within their immediate and subsequent neighborhoods until it loses interest or reaches all nodes.
-Each shared post is assumed to convey the opinion of the node on the selected topic, e.g., expressed in text as feedback to a shared news article, thereby triggering cognitive responses from neighbors.
-The reception of a post from a neighbor, triggering an opinion update on the receiving node.
-Furthermore, building upon our previous work [14], we analyze the impact of the cascade on polarization, evaluating the polarization values on the opinion vectors that maximize polarization of the classical FJ model.
-We tested the model over synthetic and real social networks.
-Our research reveals that the cascade-driven FJ model amplifies the influence of network topology compared to the classical FJ model.
-This empowers central nodes to act as opinion leaders, potentially amplifying or mitigating network polarization.
-This result implies that sharing mechanisms on online social networks can introduce imbalances in the opinion dynamics process compared to offline socialization, possibly resulting in higher levels of network polarization.
-These findings underscore the potential for online social networks to drive change and shape our social interactions.
-Finally, we validated our model with a dataset of real cascade on the Twitter platform and proved that the model is able to predict very well the real opinion dynamics.
-With our model, we can better understand the dynamics of online social networks and leverage this knowledge to create more effective communication strategies for businesses, organizations, and individuals alike.
+To address this issue, we propose the Friedkin-Johnsen on Cascade (FJC), a novel model that integrates the FJ opinion dynamics framework with information cascades.
+Specifically, we hypothesize that opinions on a selected topic are expressed and disseminated through the same flow that governs the resharing of online posts.
+A post is initially shared by certain sources and then propagates as a cascade through successive neighborhoods until it loses relevance or reaches saturation.
+When a node reshapes a post (e.g., via a retweet with a comment or framing), it not only amplifies the content but also attaches its own expressed opinion on the topic—such as a textual remark, endorsement, or critique—thereby exposing neighbors to both the content and the sharer’s stance.
+Receiving such a reposted message from a neighbor can trigger an opinion update in the recipient, linking the micro-level cascade of reposting to the macro-level evolution of opinions.
+First, we validated our FJC model on a real dataset of information cascades from Twitter, demonstrating its strong predictive power for online opinion dynamics.
+Building upon our previous work [7], we then analyzed the impact of cascade mechanisms on polarization, evaluating polarization levels using opinion vectors that maximize polarization in the classical FJ model.
+To this end, we tested the model on both synthetic and real social networks.
+Our main findings can be summarized as follows:
+  • Highly central nodes emerge as opinion leaders with disproportionate influence, capable of either intensifying or mitigating polarization.
+Notably, regardless of the extent of information spread, the sequential nature of cascades causes central nodes to encounter content early, limiting their exposure to diverse opinions and making them less susceptible to external influence.
+  • The FJC dynamics weaken the influence of peripheral nodes—even when they are numerous, stubborn, and aligned—since they often fail to reach and sway the central nodes.
+  • Information cascades can substantially shape polarization outcomes, either amplifying or reducing them, depending on how opinions spread through the network.
+These findings underscore how sharing mechanisms on online platforms introduce structural and temporal asymmetries that reshape opinion evolution and amplify the role of network topology compared to the classical FJ model.
+Furthermore, they highlight how online dynamics can exacerbate polarization beyond what is typically observed in offline social contexts.
+Neglecting the role of diffusion dynamics can therefore lead to serious misjudgments in assessing both opinion formation and polarization.
+These insights highlight the critical role of online information diffusion in shaping collective attitudes and social interactions and offer guidance for communication strategies in business, public discourse, and policy-making.
 II Related Works
-The modeling of opinion dynamics has garnered the interest of researchers for decades, with this interest further amplifying with the emergence of online social networks.
+The modeling of opinion dynamics has garnered the interest of researchers for decades, with this interest further amplified with the emergence of online social networks.
 There are several classes of opinion dynamics models, and the literature on the topic is extensive [19].
@@
 Among these models, the FJ model [3] is one of the most studied [20, 21, 22, 23, 24, 25, 26, 10, 27, 7].
-Its inner workings are explained in Seciton III-C.
+Its inner workings are explained in Section III-C.
 Recently, research on the FJ model has taken two main directions.
@@
 in [33] extended the FJ to reproduce filter bubbles.
-Differently from these works, we want to investigate the effect of the classical FJ model in a static network when altering the social debate, which is regulated by the posting flow of news on online social networks.
+Differently from these works, we investigate the effect of the classical FJ model on a static network when the social debate is shaped by the temporal flow of news posts on online social platforms.
 To model the posting flow, we mainly take hints from the very popular independent-cascade model in [36] (see Section III-B for a description of how it works), which is able to reproduce realistic diffusion dynamics [37].
-This seminal work has inspired many following studies and extentions [37, 38, 39, 40], which extend the classical model to include collaborative and/or antagonistic relationships among users or dynamic networks.
-The main goal of information diffusion is to study the rate of diffusion of viral contents depending on the parameters.
-Instead, we are not interested in the popularity of contents, but rather in their diffusion process.
+This seminal work has inspired many following studies and extensions [38, 39, 40, 41], which extend the classical model to include collaborative and/or antagonistic relationships among users, as well as dynamic networks.
+The main goal of information diffusion is to study the rate of diffusion of viral content depending on the parameters.
+Instead, we are not interested in the popularity of content, but rather in its diffusion process.
 Hence, we used the original independent-cascade model and analyzed its impact on the opinion formation process.
 Similarly to opinion dynamics, the literature on information diffusion is extensive, so we refer the interested reader to dedicated surveys for more details [42].
-The work closer to ours is [42] which proposes a model to estimate the inter-dependencies of viral contents and opinion formation.
-This model operates with two mechanisms during each time round.The first mechanism involves opinion updating, modeled using a relaxed version of the FJ model, which cannot reproduce polarization (for further details, see [14]).
-The second one is the proper interplay between information spreading and opinion adjustments.
-In this mechanism, the content spreads and its resulting popularity acts as a probability for nodes to change their innate opinion.
-In this model, the information cascade is not linked to people’s opinions when resharing the contents, nor does the flow of content sharing affect opinion evolution.
-In contrast, we assume that sharing content reflects peoples’ opinions and that the sharing process itself impacts opinion updating.
-Our goal is to understand how the propagation flow alters the FJ and investigate the impact on polarization.
-As far as we know, our work is the first approach to investigate this problem.
+The works most closely related to ours are [43, 44], which also investigate the interplay between information diffusion and opinion formation.
+While [43] adopts the Friedkin–Johnsen (FJ) model and [44] employs a bounded confidence framework, both approaches differ fundamentally from ours.
+In [43], content influences opinions only once, by shifting initial biases based on content popularity; subsequent opinion dynamics unfold independently of further information diffusion.
+In [44], opinion adoption is driven solely by similarity to the root node’s opinion and is primarily used to model content virality, not the cognitive process of opinion change.
+In contrast, our FJC model assumes that every act of content sharing conveys the sharer’s current opinion, which in turn directly influences the recipient’s belief.
+This enables us to explicitly capture how the temporal structure of cascades shapes long-term opinion dynamics and polarization—an aspect not addressed in prior work.
 III Preliminaries
@@
 We can describe an online social network platform as a graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$, where the $v$ nodes $\mathcal{V}$ represent users and the edges $\mathcal{E}$ represent social connections between them.
-In the most general form, the graph $G$ is directed and the edges $(i,j)$ represent a follow relationships.
+In the most general form, the graph $G$ is directed and the edges $(i,j)$ represent a follow relationship.
 Specifically, $(i,j)\in\mathcal{E}$ means that $i$ follows $j$, thus $j$ influences $i$ as its activity appears in $i$’s news feed.
 The set of nodes that has an influence on $i$ is thus composed of $i$’s out-neighbours, which we will denote with $\mathcal{N}^{+}(i)$.
-Instead, the set of nodes that is influenced by node $j$ is the in-neighbourhood of $j$, which we indicate with $\mathcal{N}^{-}(i)$ [^1] Please note that, generally, the neighbourhood of a node corresponds to our out-neighbourhood..
+Instead, the set of nodes that is influenced by node $j$ is the in-neighbourhood of $j$, which we indicate with $\mathcal{N}^{-}(j)$ [^1] Please note that, generally, the neighbourhood of a node corresponds to our out-neighbourhood..
 It is worth noting that in the special case of an undirected node, $\mathcal{N}^{+}(i)$ and $\mathcal{N}^{-}(i)$ are the same.
@@
 In the independent cascade model, each node can be active or inactive.
-An initial set of active nodes $v$ have a single chance to activate their neighbours $w$, and each activation is successful with probability $p_{vw}$.
-The nodes $w$ that are successfully activated, become active at the next step and try to activate their neighbours as well.
+An initial set of active nodes $v$ has a single chance to activate their neighbours $w$, and each activation is successful with probability $p_{vw}$.
+The nodes $w$ that are successfully activated become active at the next step and try to activate their neighbours as well.
 The process continues until no other nodes can be activated.
+This framework is well-suited to modeling the propagation of shared elements in online social networks, such as tweets or posts, where each user typically has one opportunity to reshare content upon seeing it, and further reshares depend on their direct followers.
 III-C Friedkin-Johnsen opinion dynamic model
@@
 The second force is a desire to reduce the discomfort people feel when their opinions differ from those around them, as described by the cognitive dissonance theory.
-Each person $i$ has a parameter $\lambda_{i}$, called susceptibilty, that determines how much weight they give to each of these forces.
-If $\lambda_{i}$ is 0, it means the person is completely stubborn and will never change its mind.
+Each person $i$ has a parameter $\lambda_{i}$, called susceptibility, that determines how much weight they give to each of these forces.
+If $\lambda_{i}$ is 0, it means the person is completely stubborn and will never change their mind.
 If $\lambda_{i}$ is 1, it means the person is completely naive and will totally get over their initial bias.
@@
 As discussed in our previous work [7], in the literature, the term “FJ model” refers to many variants of the original model.
-The update expression for the most general variant of FJ model is given by the following equation:
+The update expression for the most general variant of the FJ model is given by the following equation:
 $x_{i}(t+1)=\lambda_{i}\sum_{j=0}^{v}w_{ij}x_{j}(t)+(1-\lambda_{i})u_{i}\qquad\forall i,$ (1)
@@
 while the rest of the nodes maintain their opinion.
+Note that the term $(1-\gamma_{ij})x_{i}(t)$ balances the current expressed opinion of node $i$ against that of its neighbor $j$, with $\gamma_{ij}$ controlling the strength of influence from $j$.
 Here matrices $H=\{h_{ij}\}$ and $\Gamma=\{\gamma_{ij}\}$ replace $\Lambda$ and $W$ of the synchronous model, and are different from them.
@@
 $P_{4}(\bm{x})$ $=\sum_{i\in\mathcal{V}}|x_{i}|=\|\bm{x}\|_{1}$ (13)
-In the context of measuring polarization in social networks, the metrics $NDI$, $GDI$, and $P_{1}$ quantify polarization based on the degree of disagreement among individuals’ opinions.
+In the context of measuring polarization in social networks, the metrics $NDI$, GDIGDI, and $P_{1}$ quantify polarization based on the degree of disagreement among individuals’ opinions.
 On the other hand, metrics $P_{2}$, $P_{3}$, and $P_{4}$ consider polarization to occur when opinions are significantly distant from a neutral opinion, which is assigned a value of 0.
@@
 is referred to as the polarization value and, by definition, is positive if and only if the model is polarizing with respect to the metric $\Phi$.
-Our previous research has demonstrated that the FJ model is in fact polarizing, and a formula for obtaining polarizing initial opinion vectors for the metrics $P_{2}$, $P_{3}$, and $P_{4}$ have been identified.
-For what concerns vectors that maximise $P_{2},P_{3}$ polarization, they are correlated to the spectral properties of matrix $H=(I-\Lambda W)^{-1}(I-\Lambda)$ which transforms the initial opinions $\bm{u}$ into the final opinions $\bm{z}$ as described in Equation (2).
-In particular, we identified $\bm{u}_{B_{2}(1)}$ maximises polarization on the $L_{2}$-ball of radius 1 and $\bm{u}_{B_{2}(t)}=t\cdot\bm{u}_{B_{2}(1)}$ is its scalar multiple that maximises polarization on the $L_{2}$-ball of radius $t$.
-Furthermore, $\bm{u}_{V_{>1}}$ maximises polarization on the subspace of the specific $H$’s eigenvectors and $\bm{u}_{\max_{P_{2,3}}}$ on the entire domain.
-For what concerns vectors that maximise $P_{4}$ polarization, we identified $\bm{u}_{B_{1}(1)}$ that maximises polarization on the $L_{1}$-ball of radius 1 and $\bm{u}_{\max_{P_{4}}}$ on the entire domain.
-Unfortunately, while vectors $\bm{u}_{B_{2}(1)}$, $\bm{u}_{B_{2}(t)}$ and $\bm{u}_{B_{1}(1)}$ can be always found, $\bm{u}_{V_{>1}}$, $\bm{u}_{\max_{P_{2,3}}}$ and $\bm{u}_{\max_{P_{4}}}$ are often impossible to be computed since they are the solutions of NP-hard problems.
-However, $\bm{u}_{V_{>1}}$ can often be estimated with a heuristic.
-Please refer to [14] for further details.
+Our previous research [7] showed that the FJ model is inherently polarizing and provided formulas to derive polarizing vectors for the polarization metrics $P_{2}$, $P_{3}$, and $P_{4}$.
+These vectors are closely linked to the spectral properties of the matrix $H=(I-\Lambda W)^{-1}(I-\Lambda)$, which maps initial opinions $\bm{u}$ to final opinions $\bm{z}$ as in Equation (2).
+Depending on the considered sub-domain of the opinion space, different polarizing vectors can be identified.
+In this paper, for $P_{4}$ polarization we use $\bm{u}_{B_{1}(1)}$, while for $P_{2}$ and $P_{3}$ polarization we rely on $\bm{u}_{B_{2}(1)}$, $\bm{u}_{B_{2}(t)}$, and the heuristic $\bm{u}_{V_{>1}}^{heu}$, listed in order of increasing polarization strength.
+For a detailed mathematical treatment and derivations, we refer the reader to [7].
 In Section V we will examine the impact of information cascades when the opinion dynamic process starts with users featuring these polarizing opinions.
@@
 The purple arrows and boxes indicate the actions made within the corresponding states.
-A post may become obsolete and lose its relevance over time.
-When this occurs, nodes cease to propagate it, even if some nodes have not been reached.
-We capture this effect by defining a threshold parameter $L$, known as the interest threshold.
-Specifically, $L$ represents the maximum hop distance that a post can travel from its source to its destinations.
-Thus, once a post has reached $L$ hops, its propagation ceases.
-It is important to note that if this threshold is greater than or equal to the diameter of the network (defined as the longest shortest path between any two nodes in the graph $\mathcal{G}$), all posts will always propagate throughout the entire network, rendering the parameter $L$ ineffective in influencing the resulting cascade.
 In our model, an individual identified as the source disseminates a piece of news to the online social network.
@@
 Following this initial encounter, we maintain that nodes typically become disinterested in posts shared by others.
-Thus, we will refer to the $l$-th in-neighbourhood as the $l$-th layer or layer $l$ and we will indicate it with $\mathcal{O}_{l}^{(r)}$.
-Keeping in mind that $L$ is the interest threshold, the entire tree that comprehends all nodes in the aware state of the cascade is denoted as the sequence $\mathcal{O}^{(r)}=\left(\mathcal{O}_{l}^{(r)}\right)_{l=0}^{L}$ of all layers.
+Thus, we refer to the $l$-th in-neighbourhood as the $l$-th layer (or layer $l$), and denote it by $\mathcal{O}_{l}^{(r)}$.
+The depth of the graph is given by the eccentricity of the root node, denoted by $\epsilon^{(r)}$, which corresponds to the maximum distance from $r$ to any other node.
+Accordingly, the entire tree comprising all nodes in the aware state of the cascade is represented as the sequence $\mathcal{O}^{(r)}=\left(\mathcal{O}l^{(r)}\right)_{l=0}^{\epsilon^{(r)}}$ of all layers.
 A post $p$ generated at node $r$ spreads into the network over the tree $\mathcal{O}^{(r)}$.
 However, not all nodes in this tree receive the content, because there is a probability $\theta$ that each node will not reshare the post.
-The subtree of $\mathcal{O}^{(r)}$ that includes only nodes that have received the post is denoted as $\mathcal{T}^{(r,\theta)}=(\mathcal{T}^{(r,\theta)}_{l})_{l=0,\dots,L}$, and it holds that $\mathcal{T}^{(r,\theta)}_{l}\subseteq\mathcal{O}^{(r)}_{l}$ for each $l$.
+The subtree of $\mathcal{O}^{(r)}$ that includes only nodes that have received the post is denoted as $\mathcal{T}^{(r,\theta)}=(\mathcal{T}^{(r,\theta)}_{l})_{l=0,\dots,\epsilon^{(r)}}$, and it holds that $\mathcal{T}^{(r,\theta)}_{l}\subseteq\mathcal{O}^{(r)}_{l}$ for each $l$.
 $\mathcal{T}^{(r,\theta)}$ is a stochastic realization of the tree $\mathcal{O}^{(r)}$, as indicated by the superscript $\theta$.
@@
 Figure 2: A toy example of a network in (a) and the corresponding propagation tree for $r=1$ in (b).
-Nodes $2$ and $3$ update their opinion deterministically, while the others, do it with probability $\theta$.
+Nodes 22 and $3$ update their opinion deterministically, while the others, do it with probability $\theta$.
 We can describe it in formal terms in the following way.
 We refer to the cascade generated by the post $p$ originating in node $r$ as the collection of nodes in the tree $\mathcal{T}^{(r,\theta)}$ and we will denote it with $\mathcal{C}^{(r)}$.
-Please observe that nodes in layer 1 always belong to the cascade because they receive the post directly from the source, while those of layer $l\geq 2$ belong to the cascade if and only if they receive the post from nodes in layer $l-1$.
+Please observe that nodes in layer 1 always belong to the cascade because they receive the post directly from the source, while those of layer $l\geq 2$ belong to the cascade if and only if they receive the post from nodes in layer l1l-1.
 Let us indicate with $S_{i}$ the random variable modelling the event of resharing a post for a node $i\neq r$ (it takes the value 1 when the post is reshared, 0 otherwise).
 Then, the tree $\mathcal{T}^{(r,\theta)}$ is described as:
-$\left\{\begin{array}[]{l}\mathcal{T}^{(r,\theta)}_{0}=\{r\}\\ \mathcal{T}^{(r,\theta)}_{1}=\mathcal{N}^{-}(r)\\ \mathcal{T}^{(r,\theta)}_{l+1}=\left(\bigcup\limits_{\{i\in\mathcal{T}^{(r,\theta)}_{l}:S_{i}=1\}}\mathcal{N}^{-}(i)\right)\setminus\mathcal{T}^{(r,\theta))}_{l}.\end{array}\right.$ (15)
+[…]
 Please, remember that $\mathcal{N}^{-}(i)$ refers to the in-neighbours of node $i$.
-For our convenience, for each node $i\in\mathcal{O}^{(r)}_{l}$ we denote with $\varphi(i)$ the set of nodes who actually influence $i$’s opinion in the stochastic realization of the cascade, i.e the nodes in layer $l-1$ that are in $i$’s out-neighbour and reshare the post.
+For our convenience, for each node $i\in\mathcal{O}^{(r)}_{l}$ we denote with $\varphi(i)$ the set of nodes who actually influence $i$’s opinion in the stochastic realization of the cascade, i.e the nodes in layer l1l-1 that are in $i$’s out-neighbour and reshare the post.
 In other terms:
@@
 We will now discuss how to integrate the independent cascade model described in Section IV-A with the FJ model from Section III-C, resulting in a FJ variant (FJC) where opinions are influenced by the sharing of posts.
-In this model, opinions are influenced by the sharing of posts among nodes.
 The FJC model operates iteratively, following specific states and transitions for each node (see Figure 1).
@@
 First, these exposed nodes transition to the ”aware” state and update their opinions regarding the post.
-Following this, with a probability $\theta_{s,i}$, these nodes enter the spreader state, sharing the post again if they belong to a layer $l$ that is smaller than the maximum layer $L$.
+Following this, with a probability $\theta_{s,i}$, these nodes enter the spreader state, sharing the post again if they belong to a layer $l$ that is smaller than the maximum layer $\epsilon^{(r)}$.
 If a node does not meet this condition, the propagation of the post stops, and a new cascade is initiated.
@@
 While this assumption is stringent, it aids in designing the model .In future work, we will explore ways to relax this assumption and understand its impact on the model.
-However, in Section V-C, we demonstrate that the model is robust in real-world settings.
+In Section V-B, we discuss the model effectiveness in real-world settings.
 Now, we want to discuss the details of the opinion updating operation.
@@
 where $h_{i}$ and $\gamma_{i}$ are defined as in Equations 4 and 5, and $\varphi(i)$ is defined in the Equation (16).
-The above update rule factors in the average total influence on node $i$ by all its in-neighbours that have reshared the post, hence shared their opinion with $i$.
+The above update rule factors in the average total influence on node $i$ by all its in-neighbours that have reshared the post in the previous layer.
+This means that the average is calculated only from those neighbors who have already engaged with the post and shared their opinions, thereby influencing the opinion of node $i$.
 Note that, differently from the asynchrounous FJ update in Equation (3), node $i$ assesses the opinions of its predecessors all at once, not in a pairwise fashion.
@@
 Algorithm 1 FJC
-1:$\bm{R}$ (vector of post seeds), $\bm{u}$ (initial opinion vector)
-2:$\bm{x}$ (final opinion vector)
-3:$\bm{x}\leftarrow\bm{u}$triangleright initialize the opinion vector
+[…]
+[…]
+[…]
 4:for $r$ in $\bm{R}$ do
-5:     for $l$ in $\{1,\dots,L\}$ do $\triangleright$ iterate over layers
+5:  for $l$ in $\{1,\dots,\epsilon^{(r)}\}$ do $\triangleright$ iterate over layers
 6:   for $i$ in $\mathcal{T}^{(r,\theta)}_{l}$ do
@@
 To this end, we conducted two distinct sets of experiments.
-The first set (Sec.
-V-B) aimed to compare the opinion dynamics produced by the FJC model with those of the classical FJ model.
-For this purpose, we used three social networks—two real-world and one synthetic—and generated synthetic cascades on top of them.
-Specifically, we selected 15 seed nodes per graph to initiate the simulated cascades.
-The second set of experiments (Sec.
-V-C) focused on validating the FJC model using real-world cascade data.
+The first set of experiments (Sec.
+V-B) focused on validating the FJC model using real-world cascade data.
 Here, we evaluated how well the model can approximate opinion evolution based on traces from an online social network, where assumptions (A2) and (A3) do not always hold.
@@
 In the following, we provide a detailed description of the experimental settings and discuss the results obtained.
+The second set (Sec.
+V-C) aimed to compare the opinion dynamics produced by the FJC model with those of the classical FJ model.
+For this purpose, we used three social networks—two real-world and one synthetic—and generated synthetic cascades on top of them.
+Specifically, we selected 15 seed nodes per graph to initiate the simulated cascades.
 V-A Common experimental settings
@@
 Finally, since our primary goal is to study the impact of cascades on polarization, we initialized opinions using the polarizing vectors defined in Section III-D, ensuring a consistent and controlled setup for measuring polarization dynamics.
-V-B FJC vs FJ – with synthetic cascades
+Final opinions under cascades were then compared with those of the FJ model, whose outcomes are analytically derived via Equation (2) using the same parameters, except for the reposting probability $\theta$, which is not applicable to FJ.
+V-B FJC with real cascades
 V-B1 Dataset description and experimental setting
-In this section, we evaluate the performance of the FJC model through simulations on three social networks: the Karate Club graph [44], a Facebook graph [45], and a synthetic Barabási-Albert network [46].
-The Karate Club graph is a well-known small network with 34 nodes, while the Facebook graph includes 1519 nodes from the largest connected component of a 4039-node snapshot.
-The Barabási-Albert network, which captures the scale-free structure of real social networks, was generated with 100 nodes and parameter $k=6$, meaning each new node connects to 6 existing nodes.
-Figure 3 shows visualizations of all three graphs.
-Figure 3: Networks used for the performance evaluation. In (a) and (b) are two real social networks; in (c) a synthetic network.
-Since cascades are simulated, we randomly selected 15 seed nodes per network to initiate the posts.
-We set the resharing probability $\theta_{s,i}$ uniformly to a single value $\theta$ for all nodes, testing three scenarios: low diffusion ($\theta=0.01$), moderate diffusion ($\theta=0.8$), and maximum diffusion ($\theta=1$), where nodes always reshare posts.
-For each combination of network, susceptibility setting, reposting probability $\theta$, and initial opinion values, we performed 1000 simulation runs.
-V-B2 Opinion evolution under FJC vs FJ
-We begin by analyzing how cascades influence final opinions under the FJC model compared to the classical FJ model.
-In Figure 4, each arrow represents a node’s opinion shift—from its initial opinion (prejudice) to its final FJC opinion.
-The black dot indicates the corresponding final opinion under the FJ model.
-Panels A and B show cases where susceptibility is proportional or inversely proportional to node centrality, respectively; panel C shows uniform susceptibility ($\lambda=0.6$).
-Due to space constraints, we report results for $\theta=0.01$ and $\theta=0.8$ (shown in subpanels (i) and (ii)), and for the polarizing vectors $\bm{u}_{B_{1}(1)}$, $\bm{u}_{B_{2}(1)}$, $\bm{u}_{B_{2}(t)}$ and $\bm{u}_{V_{>1}}$, and $\bm{u}{V_{>1}}$.
-The full results are available in the SI Appendix.
-We start with panels C of Figure 4, where all nodes have equal susceptibility.
-Compared to the FJ model, highly central nodes in the FJC model tend to retain their initial opinions more strongly—visible from the shorter arrows relative to the distance to the FJ outcomes (black dots)—especially for $\theta=0.01$ (Ci) and $\theta=0.8$ (Cii).
-This leads to a wider spread of final opinions among central nodes.
-In contrast, less central nodes are more influenced and their opinions shift closer to those of central ones.
-Overall, cascades amplify the influence of network topology: central nodes act as opinion anchors, while peripheral nodes become more susceptible.
-The reposting probability $\theta$ modulates this effect—low values ($\theta=0.01$) limit opinion spread, while higher values ($\theta=0.8$) increase it.
-When susceptibility is inversely proportional to centrality (panels B), central nodes are already stubborn by design in both models, leading to similar outcomes.
-Nonetheless, cascades still shift the opinions of less central nodes more in the FJC model, as they become increasingly influenced by central ones.
-Finally, with susceptibility proportional to centrality (panels A), network topology and susceptibility act in opposing directions.
-Peripheral nodes are nearly immobile due to low susceptibility, while central nodes—despite being more susceptible—change little due to their structural position.
-As a result, final opinions under FJC remain close to initial ones, more so than in the FJ model.
-This highlights how, in cascade-driven dynamics, centrality outweighs susceptibility, with central nodes maintaining influence regardless of the surrounding opinion mass.
-Figure 4: Polarizing opinions vs.
-node centrality.
-Each arrow represents a node, starting at its initial opinion (prejudice) and ending at its final opinion under the FJC model.
-Arrow color indicates node susceptibility; black dots show the final opinions under the FJ model.
-(A) $\lambda_{i}\propto\mathrm{P}_{i}$; (B) $\lambda_{i}\propto\mathrm{P}_{i}^{-1}$; (C) $\lambda_{i}=0.6$; (i) $\theta=0.01$; (ii) $\theta=0.8$.
-In each panel, left plots correspond to $P_{4}$-polarization vectors, right plots to $P_{2}$/$P_{3}$ vectors.
-Dotted and dashed lines show the average initial and final FJC opinions.
-V-B3 Polarization
-Figure 5 shows how opinion dynamics under FJC affect polarization compared to the FJ model, measured as the change in polarization (Equation (14)) between final and initial opinions.
-We report results for each susceptibility setting—proportional (A), inversely proportional (B), and uniform (C)—and for each $\theta$ value (differently colored bars for FJC).
-Due to space constraints, we show only the polarization metric maximized by each polarizing vector in the FJ model: $P_{4}$ for $\bm{u}_{B_{1}(1)}$ and $P_{2}$ for the others.
-Full results are in the SI.
-We first consider FJC with $\theta=0.8$ and $\theta=1$ (blue and purple bars).
-When susceptibility is uniform (panel C), FJC consistently yields higher polarization than FJ, as central nodes remain influential and pull others toward their opinions.
-A similar pattern, though less pronounced, appears with inverse-centrality susceptibility (panel B), since less central nodes become more vulnerable under FJC.
-Interestingly, when susceptibility is proportional to centrality (panel A), FJC often reduces polarization compared to FJ.
-Here, central nodes are nearly immovable, and peripheral nodes, unable to influence them, shift closer to reduce cognitive dissonance—flattening the opinion landscape.
-An exception occurs in the Facebook network with $\bm{u}_{B_{1}(1)}$, where the initial opinion aligns with a central node that remains firm, attracting others and increasing polarization.
-For $\theta=0.01$ (green FJC bars), the cascade effect is weak due to limited reposting, resulting in polarization trends similar to those at higher $\theta$ but with lower magnitude.
-In some cases, the effect even reverses, leading to depolarization—primarily driven by the few nodes directly connected to the sources, which are the only ones likely to update their opinions.
-For instance, in the Karate network with proportional susceptibility and the vector $\bm{u}_{B_{1}(1)}$, most neighbors of the sources have low susceptibility and thus remain unchanged.
-The only significant shift comes from the source node itself, which lowers its opinion from 1, reducing overall polarization as measured by $P_{4}$.
-Figure 5: Polarization of FJ and FJC (with different $\theta$ values) for (A) $\lambda_{i}\sim\mathrm{P}_{i}$, (B) $\lambda_{i}\sim\mathrm{P}_{i}^{-1}$, (C) $\lambda_{i}=0.6$.
-V-C FJC with real cascades
-V-C1 Dataset description and experimental setting
 We validate the model using a real-world dataset from Bhowmick et al.
@@
 Unlike the synthetic networks in Sec.
-V-B, this is a directed graph, with multiple connected components.
+V-C, this is a directed graph, with multiple connected components.
 Since FJ-based models operate within a connected structure, we extract the largest strongly connected component (LSCC).
@@
 Initial opinions are set using the polarizing vectors described in Section III-D.
-However, due to the large size of the network, even optimized methods could not compute these vectors over the full LSCC.
-We therefore define two subgraphs:
-  • Nepal cut degree: includes only nodes with degree in the range $[150,400]$ (where the node degree spans from 2 to 53718 in the original LSCC), balancing network activity and computational feasibility.
-We extract the LSCC and retain only the cascades involving its nodes.
-  • Nepal top cascades: constructed from nodes involved in the 500 longest cascades using a heuristic for the directed Steiner Tree problem [48] to ensure that all the nodes involved in these cascades are in the subgraph.
-Table I summarizes the two subgraphs.
-The Nepal cut degree graph has more nodes but fewer edges, while Nepal top cascades is denser and includes more cascades (hence making it more likely that assumption (A3) is violated).
-TABLE I: Summary of the two real-world Nepal datasets subgraphs used for validation
+However, their computation requires to compute of the eigensystem of an $n\times n$ dense matrix, which becomes impractical due to the extensive size of the network, even when employing optimized algorithms.
+To tackle this challenge, we derived a subgraph known as Nepal top cascades, which includes nodes from the 500 longest cascading events.
+The rationale behind focusing on the longest cascades is that they inherently present a greater challenge for our model, as they involve more complex and prolonged interactions among nodes and increase the likelihood that cascades can overlap (thus, violating assumption (A3)).
+Specifically, we ensured that all nodes participating in these cascades are included in the subgraph using a heuristic for the directed Steiner Tree problem [47].
+Table I summarizes the key properties of this subgraph, which encompasses over 6,000 cascades.
+TABLE I: Summary of the real-world Nepal dataset subgraphs used for validation
 #nodes #edges #cascades max cascade length
-Nepal cut degree 30662 737541 4111 4
 Nepal top cascades 26321 4222064 6083 21
-Each dataset $D$ provides a sequence of tuples $\mathcal{C}_{D}=\big{(}(t_{1},p_{1},i_{1}),(t_{2},p_{2},i_{2}),\dots\big{)}$, where node $i_{k}$ shares post $p_{k}$ at time $t_{k}$, representing the temporal evolution of cascades.
+The dataset $D$ provides a sequence of tuples $\mathcal{C}_{D}=\big{(}(t_{1},p_{1},i_{1}),(t_{2},p_{2},i_{2}),\dots\big{)}$, where node $i_{k}$ shares post $p_{k}$ at time $t_{k}$, representing the temporal evolution of cascades.
 We assume that this sequence of reshares drives the opinion dynamics in the real world: each reshare is treated as an expression of opinion.
@@
 Table II shows the distribution of $\theta_{i}$, which is generally low, indicating that most nodes rarely reshare posts.
+Please note that these values are consistent with the retweet probabilities in [49], remaining below 0.02.
 Unlike the real dynamics, the FJC model processes cascades in a tree manner and sequentially.
 Specifically, posts are propagated from the sources to the network layer-by-layer, in accordance with assumption (A1) and each cascade starts only after the previous one ends, in accordance with assumption (A3).
-TABLE II: Statistics of followers, followed and $\theta$ in both Nepal subgraphs
-Nepal cut degree
-Min. 1st Qu. Median Mean 3rd Qu. Max. N.A.
-Followers 1.00 8.00 19.00 24.05 34.00 182.00 -
-Followees 1.00 13.00 21.00 24.05 32.00 135.00 -
-$\theta$ 0.000 0.000 0.000 0.001 0.000 1.000 8807
+TABLE II: Statistics of followers, followed and $\theta$ in the Nepal subgraph
 Nepal top cascades
-Min. 1st Qu. Median Mean 3rd Qu. Max. N.A.
-Followers 1.0 23.0 59.0 160.4 148.0 11022.0 -
-Followees 1.0 53.0 109.0 160.4 204.0 6851.0 -
-$\theta$ 0.0000 0.0000 0.0000 0.0156 0.0000 1.0000 681
-V-C2 Opinion evolution of real opinion dynamics vs FJC and FJ
-Due to the large size of the networks, we do not visualize the polarizing vectors as in Figure 4, since the plots would be unreadable.
+Min. 1st Qu. Median 3rd Qu. Max. Mean N.A.
+Followers 1.0 23.0 59.0 148.0 11022.0 160.4 -
+Followees 1.0 53.0 109.0 204.0 6851.0 160.4 -
+$\theta$ 0.0000 0.0000 0.0000 0.0000 1.0000 0.0156 681
+V-B2 Opinion evolution of real opinion dynamics vs FJC and FJ
+Due to the large size of the networks, we cannot visualize the polarizing vectors, since the plots would be unreadable.
 Instead, our main objective with these datasets is to evaluate the robustness of the FJC model against the real opinion dynamics.
 We do so by comparing the distributions of opinion shifts—i.e., the differences between final and initial opinions—for the real evolution (labeled Real) and the predictions of FJC and FJ.
-As shown in Figure 6, FJC closely matches the real dynamics in the Nepal cut degree dataset (the two curves are nearly indistinguishable) and shows reasonable agreement in the Nepal top cascades dataset.
-This is expected, as longer cascades are more likely to violate assumption (A3), reducing FJC’s accuracy.
-Nevertheless, FJC remains surprisingly robust even in this more challenging scenario.
+As shown in Figure 3, FJC closely mirrors the real dynamics observed in the Nepal top cascades dataset, outperforming the standard FJ model.
+Please note that long cascades are more likely to violate assumption (A3), reducing FJC’s accuracy.
+Nevertheless, FJC remains surprisingly reliable even in this more challenging scenario.
 Figure 3: ECDF of the opinion shift of nodes, i.e.
-final opinion minus initial opinion, in the three different opinion dynamics.
+final opinion minus initial opinion, for the Nepal top cascades in the three different opinion dynamics.
 (A) $\lambda_{i}\sim\mathrm{P}_{i}$, (B) $\lambda_{i}\sim\mathrm{P}_{i}^{-1}$, (C) $\lambda_{i}=0.6$.
-V-C3 Polarization
-Figure 7 shows polarization outcomes across different settings.
-Both cascade-based dynamics—FJC and Real—exhibit very low or even negative polarization, in sharp contrast to the FJ model.
-This suggests that, in this case, cascades tend to reduce polarization, seemingly contradicting the findings from synthetic networks in Section V-B.
-However, this can be explained by the fact that, as shown in Table II, most nodes have $\theta$ values close to zero, indicating minimal reposting activity.
-These results are therefore consistent with our earlier observation: when $\theta$ is small, opinion dynamics are largely inactive, preventing polarization from developing.
+V-B3 Polarization
+Figure 4 illustrates the polarization outcomes across different settings.
+As expected, the Real opinion dynamics remain almost entirely null, since opinion shifts are negligible in all scenarios (Figure 3).
+Similarly, the cascade-based dynamics (FJC) exhibit very low—or even slightly negative—polarization.
+Both stand in sharp contrast to the classical FJ model, which consistently produces positive polarization with substantially higher magnitude.
+This behavior can be explained by Table II, which shows that most nodes have $\theta$ values close to zero, indicating very limited reposting activity.
+Consequently, the opinion dynamics remain largely inactive, preventing polarization from emerging.
+In Section V-C, we will further validate this hypothesis on synthetic networks, also exploring scenarios with larger $\theta$ values to assess their impact on polarization.
 Figure 4: Polarization value of the different vectors in the corresponding metrics in the three different opinion dynamics.
 (A) $\lambda_{i}\sim\mathrm{P}_{i}$, (B) $\lambda_{i}\sim\mathrm{P}_{i}^{-1}$, (C) $\lambda_{i}=0.6$.
+V-C FJC vs FJ – with synthetic cascades
+V-C1 Dataset description and experimental setting
+In this section, we evaluate the performance of the FJC model through simulations on three social networks: the Karate Club graph [50], a Facebook graph [51], and a synthetic Barabási-Albert network [52].
+The Karate Club graph is a well-known small network with 34 nodes, while the Facebook graph includes 1519 nodes from the largest connected component of a 4039-node snapshot.
+The Barabási-Albert network, which captures the scale-free structure of real social networks, was generated with 100 nodes and parameter $k=6$, meaning each new node connects to 6 existing nodes.
+Figure 5 shows visualizations of all three graphs.
+Figure 5: Networks used for the performance evaluation. In (a) and (b) are two real social networks; in (c) a synthetic network.
+Since cascades are simulated, we randomly selected 15 seed nodes per network to initiate the posts, which are kept fixed across all simulations.
+We set the resharing probability $\theta_{s,i}$ uniformly to a single value $\theta$ for all nodes, and tested three different scenarios.
+In the first diffusion scenario, we set $\theta=0.01$, which is approximately the average value found in the Nepal dataset (see Table II).
+To explore other potential values, we referenced typical ranges established in the literature.
+Hodas et al.
+in [53] estimated that the retweet probability for each tweet seen is around 0.15% to 0.4%.
+Based on this, we modelled a moderate diffusion scenario with $\theta=0.1$, and for maximum diffusion, we used $\theta=0.5$.
+For each combination of network, susceptibility setting, reposting probability $\theta$, and initial opinion values, we performed 1000 independent simulation runs.
+V-C2 Opinion evolution under FJC vs FJ
+We begin by analyzing how cascades influence final opinions under the FJC model compared to the classical FJ model.
+Due to space constraints, we will present just $\bm{u}_{B_{1}(1)}$ and $\bm{u}_{V_{>1}}^{heu}$, the best local sub-optimal polarizing vectors that maximise $P_{4}$ and $P_{2}-P_{3}$ polarization, respectively.
+Please refer to the Supplemental Information (SI) for the full results.
+In Figures 6-8, each arrow represents a node’s opinion shift—from its initial opinion (prejudice) to its final FJC opinion.
+The black dot indicates the corresponding final opinion under the FJ model.
+Panels A and B show cases where susceptibility is proportional or inversely proportional to node centrality, respectively; panel C shows uniform susceptibility ($\lambda=0.6$).
+We begin our discussion with Figure 8, which corresponds to the scenario with the highest propagation ($\theta=0.5$), as this setting highlights the effects of cascades on opinion dynamics at a larger scale.
+Afterwards, we will examine how these trends change when $\theta$ is smaller.
+Let us first consider panel C, where all nodes share the same susceptibility.
+Compared to the classical FJ model, highly central nodes in the FJC model tend to retain their initial opinions more strongly—evident from the shorter arrows relative to the distance to the FJ outcomes (black dots).
+This leads to a wider spread of final opinions among central nodes.
+In contrast, peripheral nodes are more easily influenced, and their opinions shift closer to those of the central ones.
+Overall, cascades amplify the role of network topology: central nodes act as opinion anchors, while peripheral nodes become more susceptible followers.
+When susceptibility is inversely proportional to centrality (panel B), central nodes are already stubborn by design in both models.
+Final opinions are mostly similar across all networks, although we can observe slightly smaller opinion shifts in FJC.
+This reduction is due to the parameter $\theta$, which limits opinion propagation.
+This trend is further confirmed by the figures in the SI, where we also tested even larger values for this parameter: as $\theta$ increases, peripheral nodes shift more strongly in the FJC model, becoming increasingly influenced by central nodes.
+Finally, with susceptibility proportional to centrality (panel A), network topology and susceptibility act in opposing directions.
+Peripheral nodes are nearly immobile due to low susceptibility, while central nodes—despite being more susceptible—change little due to their structural position.
+As a result, final opinions under FJC remain close to initial ones, more so than in the FJ model.
+This highlights how, in cascade-driven dynamics, centrality outweighs susceptibility, with central nodes maintaining influence regardless of the surrounding opinion mass.
+Similar behaviors, though less pronounced, are observed with smaller values of $\theta$ ($\theta=0.01,0.1$) in Figures 6 and 7.
+In general, peripheral nodes exhibit reduced opinion shifts as $\theta$ decreases, reflecting their stronger dependence on the probability of content being reshared.
+Notably, in the Karate network, results remain almost unchanged across different $\theta$ values, as its small size allows most nodes to be reached by the cascade even when propagation is limited.
+Interestingly, central nodes in the FJC model maintain stable behavior even with tighter propagation.
+This resilience is an intrinsic property of the model, stemming from the structured, sequential nature of its cascades.
+In fact, highly connected nodes are typically exposed to content early, when active sources are few, making them less susceptible to external influence.
+$\bm{\theta=0.01}$
+Figure 6: Polarizing opinions vs.
+node centrality with $\theta=0.01$.
+Each arrow represents a node, starting at its initial opinion (prejudice) and ending at its final opinion under the FJC model.
+Arrow color indicates node susceptibility; black dots show the final opinions under the FJ model.
+(A) $\lambda_{i}\propto\mathrm{P}_{i}$; (B) $\lambda_{i}\propto\mathrm{P}_{i}^{-1}$; (C) $\lambda_{i}=0.6$.
+Dot-dashed and dashed lines show the average initial and final FJC opinions.
+$\bm{\theta=0.1}$
+Figure 7: Polarizing opinions vs.
+node centrality with $\theta=0.1$.
+Each arrow represents a node, starting at its initial opinion (prejudice) and ending at its final opinion under the FJC model.
+Arrow color indicates node susceptibility; black dots show the final opinions under the FJ model.
+(A) $\lambda_{i}\propto\mathrm{P}_{i}$; (B) $\lambda_{i}\propto\mathrm{P}_{i}^{-1}$; (C) $\lambda_{i}=0.6$.
+Dot-dashed and dashed lines show the average initial and final FJC opinions.
+$\bm{\theta=0.5}$
+Figure 8: Polarizing opinions vs.
+node centrality with $\theta=0.5$.
+Each arrow represents a node, starting at its initial opinion (prejudice) and ending at its final opinion under the FJC model.
+Arrow color indicates node susceptibility; black dots show the final opinions under the FJ model.
+(A) $\lambda_{i}\propto\mathrm{P}_{i}$; (B) $\lambda_{i}\propto\mathrm{P}_{i}^{-1}$; (C) $\lambda_{i}=0.6$.
+Dot-dashed and dashed lines show the average initial and final FJC opinions.
+V-C3 Polarization
+Figure 9 shows how opinion dynamics under FJC affect polarization compared to the FJ model, measured as the change in polarization (Equation (14)) between final and initial opinions.
+We report results for each susceptibility setting—proportional (A), inversely proportional (B), and uniform (C)—and for each $\theta$ value (differently colored bars for FJC).
+We show the polarization metric maximized by each polarizing vector in the FJ model: $P_{4}$ for $\bm{u}_{B_{1}(1)}$ and $P_{2}$ for the others.
+We first consider FJC with $\theta=0.5$ (purple bars).
+When susceptibility is uniform (panel C), FJC consistently yields higher polarization than FJ, as central nodes remain influential and pull others toward their opinions.
+A similar, though less pronounced, pattern appears with inverse-centrality susceptibility (panel B), since less central nodes become more vulnerable under FJC.
+An exception arises for $\bm{u}_{{V_{>1}}^{heu}}$ in the BA network, where polarization decreases because the parameter $\theta$ limits opinion propagation, as discussed in the previous section.
+Interestingly, when susceptibility is proportional to centrality (panel A), FJC often reduces polarization compared to FJ.
+Here, central nodes are nearly immovable, and peripheral nodes, unable to influence them, shift closer to reduce cognitive dissonance—flattening the opinion landscape.
+An exception occurs in the Facebook network with $\bm{u}_{B_{1}(1)}$, where the initial opinion aligns with a central node that remains firm, attracting others and increasing polarization.
+For $\theta=0.01$ and $\theta=0.1$ (green and blue bars), the cascade effect is weak due to limited reposting, resulting in polarization trends similar to those at higher $\theta$ but with lower magnitude.
+In some cases, the effect even reverses, leading to depolarization—primarily driven by the few nodes directly connected to the sources, which are the only ones likely to update their opinions.
+For instance, in the Karate network with proportional susceptibility and the vector $\bm{u}_{B_{1}(1)}$, most neighbors of the sources have low susceptibility and thus remain unchanged.
+The only significant shift comes from the source node itself, which lowers its opinion from 1, reducing overall polarization as measured by $P_{4}$.
+Figure 9: Polarization of FJ and FJC (with different $\theta$ values) for (A) $\lambda_{i}\sim\mathrm{P}_{i}$, (B) $\lambda_{i}\sim\mathrm{P}_{i}^{-1}$, (C) $\lambda_{i}=0.6$.
 VI Conclusions
@@
 To the best of our knowledge, this is the first model that allows for the investigation of how content dissemination can affect opinion formation.
+We validated the FJC model, utilising a real cascade dataset.
+Our results demonstrate that the FJC model can reliably approximate final opinions accurately, even when many of the underlying assumptions of the model are relaxed or only partially met.
 We evaluate the model using both synthetic and real social networks and highlight that the cascades reinforce the strength of central nodes in influencing opinions.
@@
 The model builds upon certain assumptions regarding content dissemination and opinion evolution.
-To validate the FJC model, we utilized a real cascade dataset.
-Our results demonstrate that the FJC model is robust and can approximate final opinions accurately, even when many of the underlying assumptions of the model are violated.

### v3

@@
 This research underscores the importance of understanding the interplay between social dynamics and information flow in shaping public discourse in the digital age.
-Index Terms:
-Information cascades, opinion dynamics, polarization, Friedkin-Johnsen model, viral content
 I Introduction
@@
 Serving both as platforms for social interaction and as primary sources of real-time news, OSNs facilitate unprecedented levels of connectivity and information dissemination [37].
-Additionally, OSNs have emerged as prominent sources of real-time news dissemination, with a substantial proportion of the population relying on these platforms for information [1].
-Political figures, journalists, and ordinary users frequently utilize OSNs to disseminate announcements and opinions, rendering them integral to contemporary information consumption patterns.
 Moreover, media organizations have implemented strategies to encourage the sharing of their content on OSNs, facilitating the seamless integration of news into users’ social interactions.
-Therefore, during our social activities online, we can also be exposed to the news even incidentally.
-However, this convergence of socialization and news consumption on OSNs has raised concerns regarding the formation of filter bubbles and echo chambers, wherein individuals are exposed primarily to content that reinforces their existing beliefs.
-Such environments pose risks of fostering biased perspectives and perpetuating misinformation, particularly with the proliferation of fake news campaigns designed to manipulate public opinion [2].
-These dynamics underscore the complex interplay between socialization, news consumption, and information dissemination in the digital era, warranting further empirical investigation and theoretical exploration within interdisciplinary frameworks.
-Understanding the mechanisms behind these phenomena requires an integrated study of how opinions evolve through social influence and how information cascades propagate within these networks.
+Therefore, during our social activities online, we can also be exposed to the news even incidentally raising concerns about filter bubbles, echo chambers, and the rapid spread of misinformation [38].
+These phenomena highlight the need to understand how information diffusion and social influence jointly shape opinion formation in online environments.
 The dynamics of opinion formation represent a multifaceted research area situated at the intersection of various disciplines, including sociology, economics, psychology, control theory, computer science, computational social science, and physics.
 Numerous models have been proposed to describe the mechanisms underlying opinion dynamics, aiming to capture observed theoretical and empirical phenomena.
-These models often account for factors such as individuals’ inclination to reduce cognitive dissonance with close acquaintances or their resistance to relinquishing deeply held biases.
 Among the prominent models in this domain is the Friedkin-Johnsen (FJ) model [25] belonging to the category of averaging models [17, 23, 25], according to which people iteratively adjust their opinions towards the average opinion of their neighbours, with the strength of social ties weighting the influence of each neighbor’s opinion.
@@
 It has demonstrated the ability to replicate various observed mechanisms in social networks, including the phenomenon of wisdom of the crowd [16] and the emergence of polarization [11] and recently, it has also been proposed as the building block to model multilateral international negotiation processes [6].
-The Friedkin-Johnsen (FJ) model has been validated in offline settings, where opinion evolution reflects direct interpersonal interactions and deliberations across various topics [9, 10].
+The Friedkin-Johnsen model has been validated in offline settings, where opinion evolution reflects direct interpersonal interactions and deliberations across various topics [26, 24].
 However, online social environments introduce distinct mechanisms of opinion exchange: individuals are exposed to their contacts’ expressed opinions indirectly, primarily through public social media activities such as posts, comments, and shares [28, 47].
@@
 This widespread resharing suggests a natural intertwining between opinion dynamics and information cascades: as users repeatedly encounter and propagate content, microscopic acts of resharing can influence macroscopic opinion evolution.
-Despite this evident overlap, the two processes have thus far been treated separately in the literature.
+Despite this evident overlap, the two processes have largely been treated separately in the literature.
 In this work, we set out to fill this gap and establish an integrated model whereby the microscopic act of content sharing bears an effect in the macroscopic phenomenon of opinion evolution in online social networks.
-To address this issue, we propose the Friedkin-Johnsen on Cascade (FJC), a novel model that integrates the FJ opinion dynamics framework with information cascades.
-Specifically, we hypothesize that opinions on a selected topic are expressed and disseminated through the same flow that governs the resharing of online posts.
-A post is initially shared by certain sources and then propagates as a cascade through successive neighborhoods until it loses relevance or reaches saturation.
-When a node reshapes a post (e.g., via a retweet with a comment or framing), it not only amplifies the content but also attaches its own expressed opinion on the topic—such as a textual remark, endorsement, or critique—thereby exposing neighbors to both the content and the sharer’s stance.
-Receiving such a reposted message from a neighbor can trigger an opinion update in the recipient, linking the micro-level cascade of reposting to the macro-level evolution of opinions.
+To address this issue, we propose the Friedkin–Johnsen on Cascades (FJC) model, a novel framework that explicitly couples opinion dynamics with information diffusion.
+FJC departs from the classical FJ model by replacing abstract, synchronous interpersonal interactions with empirically grounded cascade-based exposure: opinions are updated only when users receive reposted content through an information cascade.
+In this setting, the social influence exerted by a neighbour is activated by the resharing process itself and is therefore constrained by the temporal order, reach, and structure of cascades.
+Concretely, cascade sources initiate posts that propagate through the network according to resharing probabilities, and each resharing event conveys the sharer’s expressed opinion on the topic.
+Opinion updates thus occur asynchronously and selectively, driven by actual exposure rather than by static network adjacency.
+This mechanism links the microscopic dynamics of content diffusion to the macroscopic evolution of opinions, yielding an opinion-formation process that is inherently shaped by cascade depth, timing, and network position.
 First, we validated our FJC model on a real dataset of information cascades from Twitter, demonstrating its strong predictive power for online opinion dynamics.
@@
 To this end, we tested the model on both synthetic and real social networks.
+Using FJC, we systematically compare opinion evolution under cascade-driven exposure with the classical FJ dynamics, both on real cascade data and on controlled synthetic networks.
 Our main findings can be summarized as follows:
@@
 II Related Works
-The modeling of opinion dynamics has garnered the interest of researchers for decades, with this interest further amplified with the emergence of online social networks.
-There are several classes of opinion dynamics models, and the literature on the topic is extensive [19].
+The modeling of opinion dynamics has attracted sustained interest for decades, an interest that has further intensified with the rise of online social networks.
+The literature is extensive and encompasses a broad spectrum of models of increasing complexity [40].
 Here, we concentrate on the class of averaging models, which posit that users adjust their opinions based on those of the people around them to mitigate social discomfort.
@@
 The second direction explores the possible extensions of the FJ model to capture diverse observed features in the online social networks [7, 9, 15, 44, 19].
-In particular, these papers incorporate both attractive and repulsive relationships between users [32, 34], even assuming that the influence of a node on another can vary according to the homophily principle [35].
-Furthermore, Bhalla et al.
-in [31] analyzes the polarization effect of FJ on time-varying networks and Chitra et al.
-in [33] extended the FJ to reproduce filter bubbles.
+Several extensions of the FJ model incorporate repulsion [9, 44], homophily [19], filter bubbles [15], and temporal networks [7].
 Differently from these works, we investigate the effect of the classical FJ model on a static network when the social debate is shaped by the temporal flow of news posts on online social platforms.
-To model the posting flow, we mainly take hints from the very popular independent-cascade model in [36] (see Section III-B for a description of how it works), which is able to reproduce realistic diffusion dynamics [37].
+Beyond FJ-specific models, similar distinctions appear in the broader opinion dynamics literature.
+On the one hand, continuous-time and kinetic frameworks [2, 56] aim to capture network evolution and large-scale interaction effects.
+In continuous-time models, the underlying social network may evolve over time, while kinetic approaches model opinion interactions in a mean-field setting, often drawing analogies with epidemic spreading and allowing agents’ susceptibility to influence to depend endogenously on their opinions.
+On the other hand, recent research has emphasized prescriptive interventions aimed at mitigating polarization in social networks.
+For instance, Liu and Wen [35] propose an agent-based framework that identifies network motifs to design targeted interventions by modifying the interaction structure.
+Within this landscape, the FJC model provides a foundational, cascade-aware description of how opinions evolve over time in online environments, which can naturally inform future intervention and application-oriented extensions.
+To model the posting flow, we draw inspiration from the independent-cascade model in [33] (see Section III-B for a description of how it works), which is able to reproduce realistic diffusion dynamics [31].
 This seminal work has inspired many following studies and extensions [5, 3, 13, 29], which extend the classical model to include collaborative and/or antagonistic relationships among users, as well as dynamic networks.
-The main goal of information diffusion is to study the rate of diffusion of viral content depending on the parameters.
-Instead, we are not interested in the popularity of content, but rather in its diffusion process.
+While information diffusion studies typically focus on content popularity, we are interested in how the diffusion process itself affects opinion formation.
 Hence, we used the original independent-cascade model and analyzed its impact on the opinion formation process.
@@
 The works most closely related to ours are [52, 12], which also investigate the interplay between information diffusion and opinion formation.
-While [43] adopts the Friedkin–Johnsen (FJ) model and [44] employs a bounded confidence framework, both approaches differ fundamentally from ours.
-In [43], content influences opinions only once, by shifting initial biases based on content popularity; subsequent opinion dynamics unfold independently of further information diffusion.
+While [52] adopts the Friedkin–Johnsen model and [12] employs a bounded confidence framework, both approaches differ fundamentally from ours.
 In [12], opinion adoption is driven solely by similarity to the root node’s opinion and is primarily used to model content virality, not the cognitive process of opinion change.
@@
 This enables us to explicitly capture how the temporal structure of cascades shapes long-term opinion dynamics and polarization—an aspect not addressed in prior work.
+Tu and Neumann [52], instead, study the interaction between viral content and opinion dynamics.
+In their model, opinion evolution and information diffusion are treated as largely separate mechanisms.
+Opinion updates are governed by a relaxed version of the FJ model, while content diffusion affects opinions only through a one-time adjustment of individuals’ innate opinions, based on the final popularity of the content.
+After this initial bias shift, opinions evolve independently of further cascade activity.
+In contrast, in our framework, content diffusion and opinion dynamics are tightly coupled: sharing content is interpreted as an explicit expression of an individual’s opinion, and each exposure to reshared content can directly trigger an opinion update.
+As a result, information cascades continuously shape the opinion formation process over time, rather than acting only through an initial perturbation.
+Complementary to our approach, Scarpa et al.
+[46] develop a Markovian agent-based framework for transient opinion dynamics in spatially structured offline settings.
+While their work emphasizes short-term and location-dependent effects, our model focuses on large-scale online cascades and emergent polarization phenomena.
 III Preliminaries
@@
 This process capitalizes on the insight that stronger relationships should carry greater weight in opinion formation.
-It’s worth noting that, with this mechanism, even in cases where the social graph exhibits symmetry, the influence matrix results asymmetric.
+It is worth noting that, with this mechanism, even in cases where the social graph exhibits symmetry, the influence matrix results asymmetric.
 III-B Information Cascade
@@
 This framework is well-suited to modeling the propagation of shared elements in online social networks, such as tweets or posts, where each user typically has one opportunity to reshare content upon seeing it, and further reshares depend on their direct followers.
-III-C Friedkin-Johnsen opinion dynamic model
+This cascade model forms the basis of the proposed FJC framework and is used extensively in Sections IV-A and IV-B.
+III-C Friedkin-Johnsen Opinion Dynamics Model
 The FJ opinion dynamic model is a mathematical model that describes how people’s opinions change over time due to their mutual influences.
-It does this by representing people’s opinions as real numbers and then modelling two opposing forces that affect how those opinions evolve.
-The first force is an initial bias that each person $i$ has, which is represented by $u_{i}$.
-The second force is a desire to reduce the discomfort people feel when their opinions differ from those around them, as described by the cognitive dissonance theory.
-Each person $i$ has a parameter $\lambda_{i}$, called susceptibility, that determines how much weight they give to each of these forces.
-If $\lambda_{i}$ is 0, it means the person is completely stubborn and will never change their mind.
-If $\lambda_{i}$ is 1, it means the person is completely naive and will totally get over their initial bias.
-Node susceptibilities are generally represented by the diagonal matrix $\Lambda$, whose $i$-th diagonal element is equal to the susceptibility $\lambda_{i}$.
 The FJ model is mathematically expressed as a simple update equation that is applied to each person’s opinion at discrete times.
@@
 The update expression for the most general variant of the FJ model is given by the following equation:
-$x_{i}(t+1)=\lambda_{i}\sum_{j=0}^{v}w_{ij}x_{j}(t)+(1-\lambda_{i})u_{i}\qquad\forall i,$ (1)
+$x_{i}(t+1)=\lambda_{i}\sum_{j\in\mathcal{V}}w_{ij}x_{j}(t)+(1-\lambda_{i})u_{i}\qquad\forall i,\vskip-5.0pt$ (1)
 where $w_{ij}$ is the $(i,j)$-th entry of matrix $W$.
@@
 where $I$ is the identity matrix.
+The final opinions computed via Theorem 1 serve as the baseline in Section V for comparison with the outcomes of the Friedkin–Johnsen model on cascades.
 This version of the FJ model is synchronous, meaning that all nodes update their opinion simultaneously.
@@
 In fact, to guarantee the convergence of the model in Equation (3) to that in Equation (1), the formulation has to take into account the probability of node $i$ being randomly selected, which, in turn, depends on the degree of nodes.
-Since high-degree nodes are involved in more edges, when selecting edges at random high-degree nodes have a higher probability of being selected.
-For this reason, $H$ and $\Gamma$ should be defined from $\Lambda$ and $W$ as follows:
+To guarantee convergence to the synchronous FJ model, $H$ and $\Gamma$ should be defined from $\Lambda$ and $W$ as follows [22]:
 $h_{i}=1-\frac{1-\lambda_{i}}{d_{i}}$ (4)
@@
 In [22], Frasca et al.
-establishes the ergodic convergence of the asynchronous FJ algorithm defined by equations (3) to the classical FJ.
+establishes the ergodic convergence of the asynchronous FJ algorithm defined by Equation (3) to the classical FJ.
 Notably, while the individual state variables generally do not converge in the pointwise sense and often exhibit persistent oscillation, their time-averaged values converge asymptotically to a stable limit, which, in our case, coincides with the stable point of the FJ.
+The FJC model is developed by building upon the asynchronous version of the FJ dynamics, as detailed in Section IV-B, which allows us to model pairwise opinion influence and embed the FJ mechanism within the cascade process.
 III-D Polarization
-In our previous work [7], we reviewed and classified the most popular polarization metrics used in the related literature.
-These metrics, given an opinion vector, measure polarization according to different semantics.
-They can be summarized in the following definition.
+In [11], the most widely used polarization metrics in the related literature are discussed and classified according to their underlying semantics.
+We summarize below a selected subset of these definitions, which are used in Section V.
 Definition 1.
 For an opinion $\bm{x}=(x_{i})\in[-1,1]^{v}$ the following polarization indices are defined:
-$NDI(\bm{x})$ $=\sum_{(i,j)\in\mathcal{E}}w_{ij}(x_{i}-x_{j})^{2}$ (8)
-$GDI(\bm{x})$ $=\sum_{i,j\in\mathcal{V}:i<j}(x_{i}-x_{j})^{2}$ (9)
-$P_{1}(\bm{x})$ $=\sum_{i\in\mathcal{V}}(x_{i}-\bar{\bm{x}})^{2}=\|\bm{x}-\bar{\bm{x}}\|_{2}^{2}$ (10)
 $P_{2}(\bm{x})$ $=\frac{1}{|\mathcal{V}|}\sum_{i\in\mathcal{V}}x_{i}^{2}=\frac{1}{|\mathcal{V}|}\|\bm{x}\|_{2}^{2}$ (8)
 $P_{3}(\bm{x})$ $=\sum_{i\in\mathcal{V}}x_{i}^{2}=\|\bm{x}\|_{2}^{2}$ (9)
-$P_{4}(\bm{x})$ $=\sum_{i\in\mathcal{V}}|x_{i}|=\|\bm{x}\|_{1}$ (13)
-In the context of measuring polarization in social networks, the metrics $NDI$, GDIGDI, and $P_{1}$ quantify polarization based on the degree of disagreement among individuals’ opinions.
-On the other hand, metrics $P_{2}$, $P_{3}$, and $P_{4}$ consider polarization to occur when opinions are significantly distant from a neutral opinion, which is assigned a value of 0.
+$P_{4}(\bm{x})$ $=\sum_{i\in\mathcal{V}}|x_{i}|=\|\bm{x}\|_{1}.$ (10)
+In the context of measuring polarization in social networks, these metrics consider polarization to occur when opinions are significantly distant from a neutral opinion, which is assigned a value of 0.
 A dynamic opinion model is considered polarizing with respect to a given metric $\Phi$ if there exists at least one initial opinion vector $\bm{u}$ such that the corresponding final opinion $\bm{z}$ exhibits greater polarization, i.e., $\Phi(\bm{z})>\Phi(\bm{u})$.
@@
 is referred to as the polarization value and, by definition, is positive if and only if the model is polarizing with respect to the metric $\Phi$.
-Our previous research [7] showed that the FJ model is inherently polarizing and provided formulas to derive polarizing vectors for the polarization metrics $P_{2}$, $P_{3}$, and $P_{4}$.
+Our previous research [11] showed that the FJ model is inherently polarizing and provided formulas to derive polarizing vectors for the polarization metrics descibed above.
 These vectors are closely linked to the spectral properties of the matrix $H=(I-\Lambda W)^{-1}(I-\Lambda)$, which maps initial opinions $\bm{u}$ to final opinions $\bm{z}$ as in Equation (2).
-Depending on the considered sub-domain of the opinion space, different polarizing vectors can be identified.
-In this paper, for $P_{4}$ polarization we use $\bm{u}_{B_{1}(1)}$, while for $P_{2}$ and $P_{3}$ polarization we rely on $\bm{u}_{B_{2}(1)}$, $\bm{u}_{B_{2}(t)}$, and the heuristic $\bm{u}_{V_{>1}}^{heu}$, listed in order of increasing polarization strength.
-For a detailed mathematical treatment and derivations, we refer the reader to [7].
-In Section V we will examine the impact of information cascades when the opinion dynamic process starts with users featuring these polarizing opinions.
-Notably, we will observe that FJC model induces significantly greater polarization than the FJ for most of the scenarios studied.
+In this paper, we use the following polarizing initial conditions:
+  • $\bm{u}_{B_{1}(1)}$: maximizes polarization under $P_{4}$ within the $L_{1}$-ball of radius 1;
+  • $\bm{u}_{B_{2}(1)}$ and $\bm{u}_{B_{2}(t)}$: maximize polarization under $P_{2}$ and $P_{3}$ within the $L_{2}$-balls of radius 1 and $t$, respectively;
+  • $\bm{u}_{V_{>1}}^{heu}$: a heuristic approximation that maximizes $P_{2}$ and $P_{3}$ over a larger subdomain of the opinion space.
+Analysis of these vectors in [11] revealed which nodes contribute most to polarization shifts in the network.
+Polarization is maximized when extreme opinions are assigned to the most influential individuals in the network, while less influential or more susceptible individuals are initialized with neutral opinions.
+Accordingly, these vectors place opinions close to the boundaries of the opinion domain (i.e., near $\pm 1$) on nodes that are either highly PageRank-central or highly stubborn, while assigning opinions close to zero to nodes that are more susceptible and/or less PageRank-central.
+The analysis in [11] confirmed that PageRank centrality is the centrality measure that better captures the influence among nodes in social networks [43].
+In Section V, we focus on these polarizing initial conditions to study the effect of information cascades under worst-case polarization scenarios for the baseline FJ model.
+While alternative initial opinion profiles would generally lead to different quantitative outcomes, these configurations are known to maximize polarization under FJ dynamics.
+While they are not guaranteed to play the same role under the FJC dynamics, they provide a principled reference point to assess how the introduction of information cascades modifies polarization.
+As we will show, information cascades tend to amplify the influence of already central individuals, who are precisely emphasized by the polarizing configurations considered here.
+For this reason, these scenarios provide a principled and informative reference point to assess how cascades can amplify or mitigate polarization.
 IV The Friedkin-Johnsen on Cascades model
-In this section, we will discuss the FJC (Friedkin-Johnsen on Cascades) model.
-The FJ model is an opinion dynamic model that explains how people form their opinions in the offline world, where face-to-face interactions significantly influence the depth of relationships and consequently the shared understanding of opinions.
-This is the unique mechanism that triggers the tension resolving process, which is the brain’s reaction to cognitive dissonance.
-Information cascades, on the other hand, describe how information is spread through online social networks, where interactions are encoded as engagement with each other’s posts.
-They characterise how a single post by a source can go viral and inundate the entire network.
-Despite belonging to two different and separate research areas, opinion dynamics and information cascades are intricately linked.
-In online social networks, opinions are expressed in a cascade manner, by subsequent exposure to other contents/opinions.
-It is reasonable to expect that this unique exposure to others’ posts influences how we update our opinions.
-The amplification of virality in social networks can disrupt the balance of real-life discussions in offline networks, thus strengthening the role of central nodes.
-This is a crucial insight that has far-reaching implications for understanding opinion formation and information dissemination in online social networks.
-IV-A Revisiting the independent cascade model
-The standard independent cascade model needs to be slightly adjusted to integrate it into the FJC framework.
-To this aim, we identify three new node states, which replace the previous ones: unaware, aware and spreader.
-A node is unaware when it has not been exposed to the post yet.
-As soon as it “sees” the post from a neighbour, the node becomes aware and then it updates its opinion as described in the following sections.
-With a probability $\theta_{s,i}$ it also reshares the post, and, when it does, the node moves into the spreader state.
-An illustration of the possible transition of states for a single node is provided in Figure 1.
+In this section, we introduce the Friedkin–Johnsen on Cascades (FJC) model, which combines opinion updating with cascade-based information diffusion.
+The FJ model describes opinion formation through interpersonal influence, while information cascades describe the propagation of content in online networks.
+The FJC model explicitly integrates these mechanisms, capturing how opinions are updated in response to exposure to posts in a cascade.
+This framework provides a basis for studying how information diffusion interacts with opinion dynamics in networked environments.
+Throughout this section, we introduce several modeling assumptions, which are discussed where they become relevant in order to facilitate understanding of the proposed framework.
+For the reader’s convenience, all assumptions are summarized in Table I.
+TABLE I: Summary of modeling assumptions in the FJC framework
+ID Assumption Description
+A1 Opinion-neutral content Shared content does not carry an intrinsic opinion; opinions are conveyed only by the individuals sharing it.
+A2 Tree-structured exposure A post propagates from the seed in a tree-like manner across successive neighbour layers, with each node influenced only upon first exposure.
+A3 Sequential cascade propagation Only one post propagates at a time; simultaneous cascades are not considered.
+IV-A Integrating Independent Cascades into the Friedkin–Johnsen Framework
+The standard IC model is adapted here to integrate opinion updating into the cascade process, yielding what we refer to as the FJC framework.
+Rather than restating the full IC mechanism, we explicitly highlight the key differences with respect to the classical model.
+First, nodes transition through three states—unaware, aware, and spreader.
+Becoming aware of a post triggers an opinion update, a feature absent in the standard IC model (see Figure 1).
+Second, information propagation is constrained to a directed tree structure rooted at the source of the post.
+Nodes are organized into successive neighbourhoods of increasing distance from the source, which we refer to as layers.
+Under this assumption, each node updates its opinion only upon its first exposure to the post, that is, when it is reached by resharing neighbours in the immediately preceding layer.
+During this exposure event, a node may aggregate influence from all such neighbours, while subsequent resharings do not induce further opinion updates.
+As a consequence, the cascade unfolds over a directed tree in which each node appears in at most one layer and is influenced only by its predecessors.
 We assume (A1) that the content being shared does not express an opinion on its own, but rather the opinion is conveyed only by the person sharing it along with a comment.
@@
 In fact, including multiple sources simply affects the first level of the information cascade tree, while the rest remains unchanged.
-We assume (A2) that once the post is shared the first time by the seed node $r$, it spreads outwards like a tree, from the original source to the first, second, third, and subsequent in-neighbours, which we will refer to as the nodes in the first, second, third, … layer.
-This assumption means that the opinions of parent nodes represent the first exposure a node has in its news feed, thereby solidifying their influence.
+We assume (A2) that once the post is shared the first time by the seed node $r$, it spreads outwards like a tree, from the original source to the first, second, third, and subsequent follower neighbours (i.e.
+in-neigbours according to notation of Section III-A), which we will refer to as the nodes in the first, second, third, … layer.
+Under this assumption, each node can appear in at most one layer of the cascade tree, corresponding to its first exposure to the post.
+While a node may be influenced by multiple predecessors in the previous layer, all such influences occur simultaneously at this first exposure.
 Following this initial encounter, we maintain that nodes typically become disinterested in posts shared by others.
@@
 Accordingly, the entire tree comprising all nodes in the aware state of the cascade is represented as the sequence $\mathcal{O}^{(r)}=\left(\mathcal{O}l^{(r)}\right)_{l=0}^{\epsilon^{(r)}}$ of all layers.
+Please note that, unlike the standard IC model, where influence attempts may occur independently along multiple incoming edges, here the cascade is explicitly represented as a directed tree rooted at the source, ensuring a unique first exposure for each node.
 A post $p$ generated at node $r$ spreads into the network over the tree $\mathcal{O}^{(r)}$.
@@
 Figure 2: A toy example of a network in (a) and the corresponding propagation tree for $r=1$ in (b).
-Nodes 22 and $3$ update their opinion deterministically, while the others, do it with probability $\theta$.
+Nodes 22 and $3$ update their opinion deterministically, while the others do it with probability $\theta$.
 We can describe it in formal terms in the following way.
@@
 Please, remember that $\mathcal{N}^{-}(i)$ refers to the in-neighbours of node $i$.
+This construction intentionally departs from the classical IC formulation to isolate the effect of single-exposure, cascade-driven opinion updates, which is essential for coupling the cascade process with the Friedkin-Johnsen dynamics.
 For our convenience, for each node $i\in\mathcal{O}^{(r)}_{l}$ we denote with $\varphi(i)$ the set of nodes who actually influence $i$’s opinion in the stochastic realization of the cascade, i.e the nodes in layer l1l-1 that are in $i$’s out-neighbour and reshare the post.
@@
 Please observe that, since we want to update opinion layer by layer, we need an asynchronous version of the FJ, similar to that of Equation (3).
-We focus on the spreading process of the $n$-th post $p_{n}$ generated at node $r_{n}$ which involves nodes in the cascade $\mathcal{C}^{(r_{n})}$.
+We focus on the spreading process of the $n$-th post $p_{n}$ generated at node $r_{n}$, which involves nodes in the cascade $\mathcal{C}^{(r_{n})}$.
 As the cascade starts, the opinions of nodes are equal to $\bm{u}=\bm{x}(0)$ if $n=1$, i.e.
@@
 $x_{i}$ $(t)=$
-$=\frac{1}{|\varphi(i)|}\Bigg{\{}\sum_{j\in\varphi(i)}h_{i}\left[(1-\gamma_{ij})x_{i}(t-1)+\right$.
-$\qquad\left.+\gamma_{ij}x_{j}(t-1)\right]+(1-h_{i})u_{i}\Bigg{\}},$ (17)
-where $h_{i}$ and $\gamma_{i}$ are defined as in Equations 4 and 5, and $\varphi(i)$ is defined in the Equation (16).
+$=\frac{1}{|\varphi(i)|}\Bigg\{\sum_{j\in\varphi(i)}h_{i}\Big[(1-\gamma_{ij})x_{i}(t-1)+\gamma_{ij}x_{j}(t-1)\Big]+$
+$\hskip 18.49988pt\hskip 18.49988pt+(1-h_{i})u_{i}\Bigg\}\vskip-5.0pt$ (14)
+where $h_{i}$ and $\gamma_{i}$ are defined as in Equations (4) and (7), and $\varphi(i)$ is defined in the Equation (13).
 The above update rule factors in the average total influence on node $i$ by all its in-neighbours that have reshared the post in the previous layer.
 This means that the average is calculated only from those neighbors who have already engaged with the post and shared their opinions, thereby influencing the opinion of node $i$.
-Note that, differently from the asynchrounous FJ update in Equation (3), node $i$ assesses the opinions of its predecessors all at once, not in a pairwise fashion.
+Note that, differently from the asynchronous FJ update in Equation (3), node $i$ assesses the opinions of its predecessors all at once, not in a pairwise fashion.
 Differently from the FJ model in Equation (1), we update opinions node by node, following the order provided by the layer of tree $\mathcal{T}^{(r_{n},\theta)}$.
@@
 As described in Section III-A, the influence matrix is derived by row-normalizing $\hat{W}$: $w_{ij}=\frac{\hat{w}{ij}}{\sum_{k}\hat{w}{ik}}$.
-Following our previous work [7], we considered three strategies for setting node susceptibility.
-The first two rely on PageRank centrality, a measure that effectively captures node influence [25].
-For each node $i$ with PageRank score $\mathrm{P}_{i}$, we assigned $\lambda_{i}$ either the rescaled value of $\mathrm{P}_{i}$ or $\mathrm{P}_{i}^{-1}$, both mapped to the interval $(0,1)$, thus giving higher susceptibility to more (or less) central nodes.
-In the third strategy, all nodes were assigned a fixed susceptibility value of 0.6.
-These variations allow us to explore how susceptibility interacts with node centrality.
+Following our previous work [11], we considered three strategies for assigning node susceptibility.
+In the first two, susceptibility is coupled to PageRank centrality, which is known to be strongly correlated with node influence in linear opinion dynamics, including the Friedkin–Johnsen model [43].
+While PageRank does not exactly coincide with the influence weights induced by FJ dynamics, it provides a well-established proxy for identifying influential nodes in directed networks.
+Specifically, for each node $i$ with PageRank score $\mathrm{P}_{i}$, we assign susceptibility $\lambda_{i}$ either proportional to $\mathrm{P}_{i}$ or to $\mathrm{P}_{i}^{-1}$, with values rescaled to the interval $(0,1)$.
+These two configurations allow us to study stressed scenarios in which susceptibility is systematically aligned or anti-aligned with node influence.
+To assess the role of this coupling, we also consider a control setting in which all nodes are assigned the same susceptibility value, $\lambda_{i}=0.6$, thereby fully decoupling susceptibility from centrality.
 Finally, since our primary goal is to study the impact of cascades on polarization, we initialized opinions using the polarizing vectors defined in Section III-D, ensuring a consistent and controlled setup for measuring polarization dynamics.
@@
 Initial opinions are set using the polarizing vectors described in Section III-D.
-However, their computation requires to compute of the eigensystem of an $n\times n$ dense matrix, which becomes impractical due to the extensive size of the network, even when employing optimized algorithms.
+However, their computation requires computing the eigensystem of an $n\times n$ dense matrix, which becomes impractical due to the extensive size of the network, even when employing optimized algorithms.
 To tackle this challenge, we derived a subgraph known as Nepal top cascades, which includes nodes from the 500 longest cascading events.
@@
 Specifically, we ensured that all nodes participating in these cascades are included in the subgraph using a heuristic for the directed Steiner Tree problem [54].
-Table I summarizes the key properties of this subgraph, which encompasses over 6,000 cascades.
-TABLE I: Summary of the real-world Nepal dataset subgraphs used for validation
+Table II summarizes the key properties of this subgraph, which encompasses over 6,000 cascades.
+To assess the robustness of our results with respect to the selection of the top-cascade subgraph, we replicated the full experimental analysis on an additional sampled subgraph; the corresponding results, which are qualitatively consistent with those reported here, are presented in the SI.
+TABLE II: Summary of the real-world Nepal dataset subgraphs used for validation
 #nodes #edges #cascades max cascade length
@@
 Inspired by weighted cascade models [33, 49], we define $\theta_{i}$ as the ratio between the number of posts reshared by node $i$ and the number of posts it has seen from its followees.
-Table II shows the distribution of $\theta_{i}$, which is generally low, indicating that most nodes rarely reshare posts.
+Table III shows the distribution of $\theta_{i}$, which is generally low, indicating that most nodes rarely reshare posts.
 Please note that these values are consistent with the retweet probabilities in [45], remaining below 0.02.
@@
 Specifically, posts are propagated from the sources to the network layer-by-layer, in accordance with assumption (A1) and each cascade starts only after the previous one ends, in accordance with assumption (A3).
-TABLE II: Statistics of followers, followed and $\theta$ in the Nepal subgraph
+TABLE III: Statistics of followers, followed and $\theta$ in the Nepal subgraph
 Nepal top cascades
@@
 We do so by comparing the distributions of opinion shifts—i.e., the differences between final and initial opinions—for the real evolution (labeled Real) and the predictions of FJC and FJ.
-As shown in Figure 3, FJC closely mirrors the real dynamics observed in the Nepal top cascades dataset, outperforming the standard FJ model.
+As shown in Figure 3, FJC closely mirrors the real dynamics observed in the Nepal top cascades dataset, with the FJC and Real curves largely overlapping, outperforming the standard FJ model.
 Please note that long cascades are more likely to violate assumption (A3), reducing FJC’s accuracy.
@@
 Figure 4 illustrates the polarization outcomes across different settings.
-As expected, the Real opinion dynamics remain almost entirely null, since opinion shifts are negligible in all scenarios (Figure 3).
+The Real opinion dynamics remain almost entirely null, since opinion shifts are negligible in all scenarios (Figure 3).
 Similarly, the cascade-based dynamics (FJC) exhibit very low—or even slightly negative—polarization.
 Both stand in sharp contrast to the classical FJ model, which consistently produces positive polarization with substantially higher magnitude.
-This behavior can be explained by Table II, which shows that most nodes have $\theta$ values close to zero, indicating very limited reposting activity.
+This behavior can be explained by Table III, which shows that most nodes have $\theta$ values close to zero, indicating very limited reposting activity.
 Consequently, the opinion dynamics remain largely inactive, preventing polarization from emerging.
 In Section V-C, we will further validate this hypothesis on synthetic networks, also exploring scenarios with larger $\theta$ values to assess their impact on polarization.
-Figure 4: Polarization value of the different vectors in the corresponding metrics in the three different opinion dynamics.
+Figure 4: Polarization value of the different vectors in the corresponding metrics in the three different opinion dynamics for Nepal top cascades.
 (A) $\lambda_{i}\sim\mathrm{P}_{i}$, (B) $\lambda_{i}\sim\mathrm{P}_{i}^{-1}$, (C) $\lambda_{i}=0.6$.
@@
 The Karate Club graph is a well-known small network with 34 nodes, while the Facebook graph includes 1519 nodes from the largest connected component of a 4039-node snapshot.
-The Barabási-Albert network, which captures the scale-free structure of real social networks, was generated with 100 nodes and parameter $k=6$, meaning each new node connects to 6 existing nodes.
-Figure 5 shows visualizations of all three graphs.
-Figure 5: Networks used for the performance evaluation. In (a) and (b) are two real social networks; in (c) a synthetic network.
-Since cascades are simulated, we randomly selected 15 seed nodes per network to initiate the posts, which are kept fixed across all simulations.
+The Barabasi-Albert network, which captures the scale-free structure of real social networks, was generated with 100 nodes and parameter $k=6$, meaning each new node connects to 6 existing nodes.
+Please refer to the SI for a graphic visualization of the graphs.
+Since cascades are simulated, we randomly selected 15 seed nodes per network to initiate the posts and kept them fixed across all simulations, in order to control for variability due to source selection.
+The properties of the selected seed nodes and their centrality profiles are analyzed in the SI, where we show that this choice does not introduce systematic bias.
 We set the resharing probability $\theta_{s,i}$ uniformly to a single value $\theta$ for all nodes, and tested three different scenarios.
-In the first diffusion scenario, we set $\theta=0.01$, which is approximately the average value found in the Nepal dataset (see Table II).
+In the first diffusion scenario, we set $\theta=0.01$, which is approximately the average value found in the Nepal dataset (see Table III).
 To explore other potential values, we referenced typical ranges established in the literature.
@@
 We begin by analyzing how cascades influence final opinions under the FJC model compared to the classical FJ model.
-Due to space constraints, we will present just $\bm{u}_{B_{1}(1)}$ and $\bm{u}_{V_{>1}}^{heu}$, the best local sub-optimal polarizing vectors that maximise $P_{4}$ and $P_{2}-P_{3}$ polarization, respectively.
-Please refer to the Supplemental Information (SI) for the full results.
-In Figures 6-8, each arrow represents a node’s opinion shift—from its initial opinion (prejudice) to its final FJC opinion.
+Due to space constraints, we will present $\bm{u}_{B_{1}(1)}$ and $\bm{u}_{V_{>1}}^{heu}$, which are the local sub-optimal polarizing vectors that maximise $P_{4}$ and $P_{2}\text{-}P_{3}$ polarization, respectively, in the largest subdomain.
+Please refer to the SI for the full results.
+In Figures 5-7, each arrow represents a node’s opinion shift—from its initial opinion (prejudice) to its final FJC opinion.
 The black dot indicates the corresponding final opinion under the FJ model.
@@
 This highlights how, in cascade-driven dynamics, centrality outweighs susceptibility, with central nodes maintaining influence regardless of the surrounding opinion mass.
-Similar behaviors, though less pronounced, are observed with smaller values of $\theta$ ($\theta=0.01,0.1$) in Figures 6 and 7.
+Similar behaviors, though less pronounced, are observed with smaller values of $\theta$ ($\theta=0.01,0.1$) in Figures 5 and 6.
 In general, peripheral nodes exhibit reduced opinion shifts as $\theta$ decreases, reflecting their stronger dependence on the probability of content being reshared.
@@
 This study explores the dynamics of opinion formation and dissemination within online social networks, particularly under the influence of information cascades.
-Specifically, the paper proposes the Friedkin Johnsen on Cascade model (FJC), which integrates opinion dynamics and information cascades.
+Specifically, the paper proposes the Friedkin-Johnsen on Cascade model (FJC), which integrates opinion dynamics and information cascades.
 To the best of our knowledge, this is the first model that allows for the investigation of how content dissemination can affect opinion formation.